### v1

# Towards Controllable Speech Synthesis in the Era of Large Language Models: A Survey

Abstract
Text-to-speech (TTS), also known as speech synthesis, is a prominent research area that aims to generate natural-sounding human speech from text.  Recently, with the increasing industrial demand, TTS technologies have evolved beyond synthesizing human-like speech to enabling controllable speech generation. This includes fine-grained control over various attributes of synthesized speech such as emotion, prosody, timbre, and duration. Besides, advancements in deep learning, such as diffusion and large language models, have significantly enhanced controllable TTS over the past several years. In this paper, we conduct a comprehensive survey of controllable TTS, covering approaches ranging from basic control techniques to methods utilizing natural language prompts, aiming to provide a clear understanding of the current state of research. We examine the general controllable TTS pipeline, challenges, model architectures, and control strategies, offering a comprehensive and clear taxonomy of existing methods. Additionally, we provide a detailed summary of datasets and evaluation metrics and shed some light on the applications and future directions of controllable TTS. To the best of our knowledge, this survey paper provides the first comprehensive review of emerging controllable TTS methods, which can serve as a beneficial resource for both academic researchers and industry practitioners.
Index Terms:
Text-to-speech, controllable TTS, speech synthesis, TTS survey, large language models, diffusion models.
I Introduction
Speech synthesis, also broadly known as text-to-speech (TTS), is a long-time developed technique that aims to synthesize human-like voices from text [1, 2], and it has extensive applications in our daily lives, such as health care [3, 4], personal assistants [5], entertainment [6, 7], and robotics [8, 9]. Recently, TTS has gained significant attention with the rise of large language model (LLM)-powered chatbots, such as ChatGPT [10] and Llama [11], due to its naturalness and convenience for human-computer interaction. Meanwhile, the ability to achieve fine-grained control over synthesized speech attributes, such as emotion, prosody, timbre, and duration, has become a hot research topic in both academia and industry, driven by its vast potential for diverse applications.
Deep learning [12] has made great progress in the past decade due to exponentially growing computational resources like GPUs [13], leading to the explosion of numerous great works on TTS [14, 15, 16, 17]. These methods can synthesize human speech with better quality [14] and can achieve fine-grained control of the generated voice [18, 19, 20, 21, 22]. Besides, some recent works synthesize speech given multi-modal input, such as face images [23, 24], cartoons [7], and videos [25]. Moreover, with the fast development of open-source LLMs [11, 26, 27, 28, 29], some researchers propose to synthesize fine-grained controllable speech with natural language description [30, 31, 32], coining a new way to generate custom speech voices. Meanwhile, powering LLMs with speech synthesis has also been a hot topic in the last few years [33, 34, 35]. In recent years, a wide range of TTS methods has emerged, making it essential for researchers to gain a comprehensive understanding of current research trends, particularly in controllable TTS, to identify promising future directions in this rapidly evolving field. Consequently, there is a pressing need for an up-to-date survey of TTS techniques. While several existing surveys address parametric-based approaches [36, 37, 38, 39, 40, 41] and deep learning-based TTS [42, 43, 44, 45, 46, 47, 48], they largely overlook the controllability of TTS. Additionally, these surveys do not cover the advancements in recent years, such as natural language description-based TTS methods.
This paper provides a comprehensive and in-depth survey of existing and emerging TTS technologies, with a particular focus on controllable TTS methods. Fig. 1 demonstrates the development of controllable TTS methods in recent years, showing their backbones, feature representations, and control abilities. The remainder of this section begins with a brief comparison between this survey and previous ones, followed by an overview of the history of controllable TTS technologies, ranging from early milestones to state-of-the-art advancements. Finally, we introduce the taxonomy and organization of this paper.
Figure 1: A summary of representative controllable TTS methods in recent years and their model architectures, feature representations, and control abilities. There are additional network structures, such as VAE-based models, that are not included in this figure. For more details, refer to Tables IV and III.
I-A Comparison with Existing Surveys
Several survey papers have reviewed TTS technologies, spanning early approaches from previous decades [36, 37, 49, 40] to more recent advancements [43, 42, 50]. However, to the best of our knowledge, this paper is the first to focus specifically on controllable TTS. The key differences between this survey and prior work are summarized as follows:
Different scope. Klatt et al. [36] provided the first comprehensive survey on formant, concatenative, and articulatory TTS methods, with a strong emphasis on text analysis. In the early 2010s, Tabet et al. [49] and King et al. [40] explored rule-based, concatenative, and HMM-based techniques. Later, the advent of deep learning catalyzed the emergence of numerous neural-based TTS methods. Therefore, Ning et al. [43] and Tan et al. [42] have conducted extensive surveys on neural-based acoustic models and vocoders, while Zhang et al. [50] presented the first review of diffusion model-based TTS techniques. However, these studies offer limited discussion on the controllability of TTS systems. To address this gap, we present the first comprehensive survey of TTS methods through the lens of controllability, providing an in-depth analysis of model architectures and strategies for controlling synthesized speech.
Close to current demand. With the rapid development of hardware (i.e., GPUs) and artificial intelligence techniques (i.e., transformers, LLMs, diffusion models) in the last few years, the demand for controllable TTS is becoming increasingly urgent due to its broad applications in industries such as filmmaking, gaming, robots, and personal assistants. Despite this growing need, existing surveys pay little attention to control methods in TTS technologies. To bridge this gap, we propose a systematic analysis of current controllable TTS methods and the associated challenges, offering a comprehensive understanding of the research state in this field.
New insights and directions. This survey offers new insights through a comprehensive analysis of model architectures and control methods in controllable TTS systems. Additionally, it provides an in-depth discussion of the challenges associated with various controllable TTS tasks. Furthermore, we address the question: “Where are we on the path to fully controllable TTS technologies?”, by examining the relationship and gap between current TTS methods and industrial requirements. Based on these analyses, we identify promising directions for future research on TTS technologies.
Table I summarizes representative surveys and this paper in terms of main focus and publication year.
TABLE I: Comparison with representative TTS surveys.
Survey Main Focus Year
Klatt et al. [36] Rule-based and concatenative TTS 1987
Tabet et al. [49] Rule-based, concatenative, and parametric TTS 2011
King et al. [40] Parametric TTS and performance measurement 2014
Tan et al. [42] Neural-based, efficient, and expressive TTS 2021
Zhang et al. [50] Diffusion-based TTS and speech enhancement 2023
Ours Controllable TTS and evaluation 2024
I-B The History of Controllable TTS
Controllable TTS aims to control various aspects of synthesized speech, such as pitch, energy, speed/duration, prosody, timbre, emotion, gender, or high-level styles. This subsection briefly reviews the history of controllable TTS ranging from early approaches to the state-of-arts in recent years.
Early approaches. Before the prevalence of deep neural networks (DNNs), controllable TTS technologies were built primarily on rule-based, concatenative, and statistical methods. These approaches enable some degree of customization and control, though they were constrained by the limitations of the underlying models and available computational resources. 1) Rule-based TTS systems [51, 52, 53, 54], such as formant synthesis, were among the earliest methods for speech generation. These systems use manually crafted rules to simulate the speech generation process by controlling acoustic parameters such as pitch, duration, and formant frequencies, allowing explicit manipulation of prosody and phonetic details through rule adjustments. 2) Concatenative TTS [55, 56, 57, 58], which dominated the field in the late 1990s and early 2000s, synthesize speech by concatenating pre-recorded speech segments, such as phonemes or diphones, stored in a large database [59]. These methods can modify the prosody by manipulating the pitch, duration, and amplitude of speech segments during concatenation. They also allow limited voice customization by selecting speech units from different speakers. 3) Parametric methods, particularly HMM-based TTS [60, 61, 62, 63, 64, 65], gained prominence in the late 2000s. These systems model the relationships between linguistic features and acoustic parameters, providing more flexibility in controlling prosody, pitch, speaking rate, and timbre by adjusting statistical parameters. Some HMM-based systems also supported speaker adaptation [66, 67] and voice conversion [68, 69], enabling voice cloning to some extent. Besides, emotion can also be limitedly controlled by some of these methods [70, 71, 60, 72]. In addition, they required less storage compared to concatenative TTS and allowed smoother transitions between speech units.
Neural-based synthesis. Neural-based TTS technologies emerged with the advent of deep learning, significantly advancing the field by enabling more flexible, natural, and expressive speech synthesis. Unlike traditional methods, neural-based TTS leverages DNNs to model complex relationships between input text and speech, facilitating nuanced control over various speech characteristics. Early neural TTS systems, such as WaveNet [73] and Tacotron [74] laid the groundwork for controllability. 1) Controlling prosody features like rhythm and intonation is vital for generating expressive and contextually appropriate speech. Neural-based TTS models achieve prosody control through explicit conditioning or learned latent representations [75, 15, 76, 77, 78]. 2) Speaker control has also gained significant improvement in neural-based TTS through speaker embeddings or adaptation techniques [79, 80, 81, 82]. 3) Besides, emotionally controllable TTS [83, 31, 20, 22, 32] has become a hot topic due to the strong modeling capability of DNNs, enabling the synthesis of speech with specific emotional tones such as happiness, sadness, anger, or neutrality. These systems go beyond producing intelligible and natural-sounding speech, focusing on generating expressive output that aligns with the intended emotional context. 4) Neural-based TTS can also manipulate timbre (vocal quality) [78, 84, 85, 14, 86, 87] and style (speech mannerisms) [88, 89, 90], allowing for creative and personalized applications. These techniques lead to one of the most popular research topics, i.e., zero-shot TTS (particularly voice cloning) [82, 78, 91, 92]. 5) Fine-grained content and linguistic control also become more powerful [93, 94, 95, 96]. These methods can emphasize or de-emphasize specific words or adjust the pronunciation of phonemes through speech editing or generation techniques.
Neural-based TTS technologies represent a significant leap in the flexibility and quality of speech synthesis. From prosody and emotion to speaker identity and style, these systems empower diverse applications in fields such as entertainment, accessibility, and human-computer interaction.
LLM-based synthesis. Here we pay special attention to LLM-based synthesis methods due to their superior context modeling capabilities compared to other neural-based TTS methods. LLMs, such as GPT [97, 98], T5 [99], and PaLM [100], have revolutionized various natural language processing (NLP) tasks with their ability to generate coherent, context-aware text. Recently, their utility has expanded into controllable TTS technologies [101, 102, 103, 104, 17]. For example, users can synthesize the target speech by describing its characteristics, such as: “A young girl says ‘I really like it, thank you!’ with a happy voice”, making speech generation significantly more intuitive and user-friendly. Specifically, an LLM can detect emotional intent in sentences (e.g., “I’m thrilled” → happiness, “This is unfortunate” → sadness). The detected emotion is encoded as an auxiliary input to the TTS model, enabling modulation of acoustic features like prosody, pitch, and energy to align with the expressed sentiment. By leveraging LLMs’ capabilities in understanding and generating rich contextual information, these systems can achieve enhanced and fine-grained control over various speech attributes such as prosody, emotion, style, and speaker characteristics [105, 31, 106]. Integrating LLMs into TTS systems represents a significant step forward, enabling more dynamic and expressive speech synthesis.
I-C Organization of This Survey
This paper first presents a comprehensive and systematic review of controllable TTS technologies, with a particular focus on model architectures, control methodologies, and feature representations. To establish a foundational understanding, this survey begins with an introduction to the TTS pipeline in Section II. While our focus remains on controllable TTS, Section III examines seminal works in uncontrollable TTS that have significantly influenced the field’s development. Section IV provides a thorough investigation into controllable TTS methods, analyzing both their model architectures and control strategies. Section V presents a comprehensive review of datasets and evaluation metrics. Section VI provides an in-depth analysis of the challenges encountered in achieving controllable TTS systems and discusses future directions. Section VII explores the broader impacts of controllable TTS technologies and identifies promising future research directions, followed by the conclusion in Section VIII.
II TTS Pipeline
Figure 2: The general pipeline of controllable TTS from the perspective of network structure. Linguistic analysis is necessary for parametric and a few neural-based methods but is no longer needed for most modern neural-based methods. In this paper, we only review neural-based controllable TTS methods and do not investigate acoustic features (e.g., MCC [107], LSP [108], F0 [109]) used in early TTS methods.
In this section, we elaborate on the general pipeline that supports controllable TTS technologies, including acoustic models, speech vocoders, and feature representations. Fig.2 depicts the general pipeline of controllable TTS, containing various model architectures and feature representations, but the control strategies will be discussed in Section IV. Readers can jump to Section III if familiar with TTS pipelines.
II-A Overview
A TTS pipeline generally contains three key components, i.e., linguistic analyzer, acoustic model, speech vocoder, and with a conditional input, e.g., prompts, for controllable speech synthesis. Besides, some end-to-end methods use a single model to encode the input and decode the speech waveforms without generating intermediate features like mel-spectrograms [110]. Linguistic analyzer aims to extract linguistic features, e.g., phoneme duration and position, syllable stress, and utterance level, from the input text, which is a necessary step in HHM-based methods [64, 65] and a few neural-based methods [111, 112], but is time-consuming and error-prone. Acoustic model is a parametric or neural model that predicts the acoustic features from the input texts. Modern neural-based acoustic models like Tacotron [74] and later works [15, 76, 113] directly take character [114] or word embeddings [115] as the input, which is much more efficient than previous methods. Speech vocoder is the last component that converts the intermediate acoustic features into a waveform that can be played back. This step bridges the gap between the acoustic features and the actual sounds produced, helping to generate high-quality, natural-sounding speech [73, 116]. Tan et al. [42] have presented a comprehensive and detailed review of acoustic models and vocoders. Therefore, the following subsections will briefly introduce some representative acoustic models and speech vocoders, followed by a discussion of acoustic feature representations.
II-B Acoustic Models
Acoustic modeling is a crucial step in TTS because it ensures the generated acoustic features capture the subtleties of human speech. By accurately modeling acoustic features, modern TTS systems can help generate high-quality and expressive audio that sounds close to human speech.
Parametric models. Early acoustic models rely on parametric approaches, where predefined rules and mathematical functions are utilized to model speech generation. These models often utilize HMMs to capture acoustic features from linguistic input and generate acoustic features by parameterizing the vocal tract and its physiological properties such as pitch and prosody [117, 118, 119, 120, 72, 71]. These methods have relatively low computational costs and can produce a range of voices by adjusting model parameters. However, the speech quality of these methods is robotic and lacks natural intonation, and the expressiveness is also limited [120, 72].
RNN-based models. Recurrent Neural Networks (RNNs) proved particularly effective in early neural-based TTS due to their ability to model sequential data and long-range dependencies, which helps in capturing the sequential nature of speech, such as the duration and natural flow of phonemes. Typically, these models have an encoder-decoder architecture, where an encoder encodes input linguistic features, such as phonemes or text, into a fixed-dimensional representation, and the decoder sequentially decodes this representation into acoustic features (e.g., mel-spectrogram frames) that capture the frequency and amplitude of sound over time. Tacotron 2 [75] is one of the pioneering TTS models that uses RNNs with an attention mechanism, which helps align the text sequence with the generated acoustic features. It takes raw characters as input and produces mel-spectrogram frames, which are subsequently converted to waveforms. Another example is MelNet [121], which leverages autoregressive modeling to generate high-quality mel-spectrograms, demonstrating versatility in generating both speech and music, achieving high fidelity and coherence across temporal scales.
CNN-based models. Unlike RNNs, which process sequential data frame by frame, CNNs process the entire sequence at once by applying filters across the input texts. This parallel approach enables faster training and inference, making CNN-based TTS particularly appealing for real-time and low-latency applications. Furthermore, by stacking multiple convolutional layers with varying kernel sizes or dilation rates, CNNs can capture both short-range and long-range dependencies, which are essential for natural-sounding speech synthesis. Deep Voice [122] is one of the first prominent CNN-based TTS models by Baidu, designed to generate mel-spectrograms directly from phoneme or character input. ParaNet [123] also utilizes a RNN model to achieve sequence-to-sequence mel-spectrogram generation. It uses a non-autoregressive architecture, which enables significantly faster inference by predicting multiple time steps simultaneously.
Transformer-based models. Transformer model [124] uses self-attention layers to capture relationships within the input sequence, making them well-suited for tasks requiring an understanding of global contexts, such as prosody and rhythm in TTS. Transformer-based TTS models often employ an encoder-decoder architecture, where the encoder processes linguistic information (e.g., phonemes or text) and captures contextual relationships, and the decoder generates acoustic features (like mel-spectrograms) from these encoded representations, later converted to waveforms by a vocoder. TransformerTTS [125] is one of the first TTS models that apply transformers to synthesize speech from text. It utilizes a standard encoder-decoder transformer architecture and relies on multi-head self-attention mechanisms to model long-term dependencies, which helps maintain consistency and natural flow in speech over long utterances. FastSpeech [15] is a non-autoregressive model designed to overcome the limitations of autoregressive transformers in TTS, achieving faster synthesis than previous methods. It introduces a length regulator to align text with output frames, enabling the control of phoneme duration. FastSpeech 2 [76] extends FastSpeech by adding pitch, duration, and energy predictors, resulting in more expressive and natural-sounding speech.
LLM-based models. LLMs [126, 97, 11, 26], known for their large-scale pre-training on text data, have shown remarkable capabilities in natural language understanding and generation. LLM-based TTS models generally use a text description to guide the mel-spectrogram generation, where the acoustic model processes the input text to generate acoustic tokens that capture linguistic and contextual information, such as tone, sentiment, and prosody. For example, PromptTTS [101] uses a textual prompt encoded by BERT [126] to guide the acoustic model on the timbre, tone, emotion, and prosody desired in the speech output. PromptTTS first generates mel-spectrograms with token embeddings and then converts them to audio using a vocoder. InstructTTS [105] generates expressive and controllable speech using natural language style prompts. It leverages discrete latent representations of speech and integrates natural language descriptions to guide the synthesis process, which bridges the gap between TTS systems and natural language interfaces, enabling fine-grained style control through intuitive prompts.
Other acoustic models. In TTS, GANs [127, 128, 129], VAEs [130, 18], and diffusion models [113, 131] can also be used as acoustic models. Flow-based methods [132, 133] are also popular in waveform generation. Refer to the survey paper from Tan et al. [42] for more details.
The choice of an acoustic model depends on the specific requirements and is a trade-off between synthesis quality, computational efficiency, and flexibility. For real-time applications, CNN-based or lightweight transformer-based models are preferable, while for high-fidelity, expressive speech synthesis, transformer-based and LLM-based models are better suited.
II-C Speech Vocoders
Vocoders are essential for converting acoustic features such as mel-spectrograms into intelligible audio waveforms and are vital in determining the naturalness and quality of synthesized speech. We broadly categorize existing vocoders according to their model architectures, i.e., RNN-, CNN-, GAN-, and diffusion-based vocoders.
RNN-based vocoders. Unlike traditional vocoders [134, 135] that depend on manually designed signal processing pipelines, RNN-based vocoders [136, 137, 138, 139] leverage the temporal modeling capabilities of RNNs to directly learn the complex patterns in speech signals, enabling the synthesis of natural-sounding waveforms with improved prosody and temporal coherence. For instance, WaveRNN [137] generates speech waveforms sample-by-sample using a single-layer recurrent neural network, typically with Gated Recurrent Units (GRU). It improves upon earlier neural vocoders like WaveNet [73] by significantly reducing the computational requirements without sacrificing audio quality. MB-WaveRNN [139] extends WaveRNN by incorporating a multi-band decomposition strategy, where the speech waveform is divided into multiple sub-bands, with each sub-band synthesized at a lower sampling rate. These sub-bands are then combined to reconstruct the full-band waveform, thereby accelerating the synthesis process while preserving audio quality.
CNN-based vocoders. By leveraging the parallel nature of convolutional operations, CNN-based vocoders [73, 140, 141] can generate high-quality speech more efficiently, making them ideal for real-time applications. A key strength of CNN-based vocoders is their ability to balance synthesis quality and efficiency. However, they often require extensive training data and careful hyperparameter tuning to achieve optimal performance. WaveNet [73] is a probabilistic autoregressive model that generates waveforms sample by sample conditioned on all preceding samples and auxiliary inputs, such as linguistic features and mel-spectrograms. It employs stacks of dilated causal convolutions, enabling long-range dependence modeling in speech signals without relying on recurrent connections. Parallel WaveNet [140] addresses WaveNet’s inference speed limitations while maintaining comparable synthesis quality. It introduces a non-autoregressive mechanism based on a teacher-student framework, where the original WaveNet (teacher) distills knowledge into a student model. The student generates samples in parallel, enabling real-time synthesis without waveform quality degradation.
GAN-based vocoders. GANs have been widely adopted in vocoders for high-quality speech generation [142, 143, 116, 144, 145], leveraging adversarial losses to improve realism. GAN-based vocoders typically consist of a generator that produces waveforms conditioned on acoustic features, such as mel-spectrograms, and a discriminator that distinguishes between real and synthesized waveforms. Models like Parallel WaveGAN [144] and HiFi-GAN [116] have demonstrated the effectiveness of GANs in vocoding by introducing tailored loss functions, such as multi-scale and multi-resolution spectrogram losses, to ensure naturalness in both time and frequency domains. These models can efficiently handle the complex, non-linear relationships inherent in speech signals, resulting in high-quality synthesis. A key advantage of GAN-based vocoders is their parallel inference capability, enabling real-time synthesis with lower computational costs compared to autoregressive models. However, training GANs can be challenging due to instability and mode collapse. Despite these challenges, GAN-based vocoders continue to advance the state-of-the-art in neural vocoding, offering a compelling combination of speed and audio quality.
Diffusion-based vocoders. Inspired by diffusion probabilistic models [146] that have shown success in visual generation tasks, diffusion-based vocoders [147, 113, 148, 149, 150] present a novel approach to natural-sounding speech synthesis. The core mechanism of diffusion-based vocoders involves two stages: a forward process and a reverse process. In the forward process, clean speech waveforms are progressively corrupted by adding noise in a controlled manner, creating a sequence of intermediate noisy representations. During training, the model learns to reverse this process, progressively denoising the corrupted signal to reconstruct the original waveform. Diffusion-based vocoders, such as WaveGrad [149] and DiffWave [148], have demonstrated remarkable performance in generating high-fidelity waveforms while maintaining temporal coherence and natural prosody. They offer advantages over previous vocoders, including robustness to over-smoothing [151] and the ability to model complex data distributions. However, their iterative sampling process can be computationally intensive, posing challenges for real-time applications.
Other vocoders. There are also many other types of vocoders such as flow-based [152, 153, 154, 155, 156] and VAE-based vocoders [157, 158, 159]. These methods provide unique strengths for speech synthesis such as efficiency and greater flexibility in modeling complex speech variations. Readers can refer to the survey paper from Tan et al. [42] for more details.
The choice of vocoder depends on various factors. While high-quality models like GAN-based and diffusion-based vocoders excel in naturalness, they may not be suitable for real-time scenarios. On the other hand, models like Parallel WaveNet [140] balance quality and efficiency for practical use cases. The best choice will ultimately depend on the specific use case, available resources, and the importance of factors such as model size, training data, and inference speed.
II-D Fully End-to-end TTS models
Fully end-to-end TTS methods [160, 161, 76, 159, 162] directly generate speech waveforms from textual input, simplifying the “acoustic model → vocoder” pipeline and achieving efficient speech generation. Char2Wav [160] is an early neural text-to-speech (TTS) system that directly synthesizes speech waveforms from character-level text input. It integrates two components and jointly trains them: a recurrent sequence-to-sequence model with attention, which predicts acoustic features (e.g., mel-spectrograms) from text, and a SampleRNN-based neural vocoder [136] that generates waveforms from these features. Similarly, FastSpeech 2s [76] directly synthesizes speech waveforms from texts by extending FastSpeech 2 [76] with a waveform decoder, achieving high-quality and low-latency synthesis.   VITS [159] is another fully end-to-end TTS framework. It integrates a variational autoencoder (VAE) with normalizing flows [163] and adversarial training, enabling the model to learn latent representations that capture the intricate variations in speech, such as prosody and style. VITS combines non-autoregressive synthesis with stochastic latent variable modeling, achieving real-time waveform generation without compromising naturalness. There are more end-to-end TTS models such as Tacotron [74], ClariNet [161], and EATS [162], refer to another survey [42] for more details. End-to-end controllable methods that emerged in recent years will be discussed in Section IV.
TABLE II: Popular open-source speech quantization methods.
SSCP: Self-supervised context prediction, RVQ: Residual vector quantization [168].
II-E Acoustic Feature Representations
In TTS, the choice of acoustic feature representations impacts the model’s flexibility, quality, expressiveness, and controllability. This subsection investigates continuous representations and discrete tokens as shown in Fig.2, along with their pros and cons for TTS applications.
Continuous representations. Continuous representations (e.g., mel-spectrograms) of intermediate acoustic features use a continuous feature space to represent speech signals. These representations often involve acoustic features that capture frequency, pitch, and other characteristics without discretizing the signal. The advantages of continuous features are: 1) Continuous representations retain fine-grained detail, enabling more expressive and natural-sounding speech synthesis. 2) Since continuous features inherently capture variations in tone, pitch, and emphasis, they are well-suited for prosody control and emotional TTS. 3) Continuous representations are more robust to information loss and can avoid quantization artifacts, allowing smoother, less distorted audio. GAN-based [116, 144, 145] and diffusion-based methods [148, 147] often utilize continuous feature representations, i.e., mel-spectrograms. However, continuous representations are typically more computationally demanding and require larger models and memory, especially in high-resolution audio synthesis.
Discrete tokens. In discrete token-based TTS, the intermediate acoustic features (e.g., quantized units or phoneme-like tokens) are discrete values, similar to words or phonemes in languages. These are often produced using quantization techniques or learned embeddings, such as HuBERT [166] and SoundStream [168]. The advantages of discrete tokens are: 1) Discrete tokens can encode phonemes or sub-word units, making them concise and less computationally demanding to handle. 2) Discrete tokens often allow TTS systems to require fewer samples to learn and generalize, as the representations are compact and simplified. 3) Using discrete tokens simplifies cross-modal TTS applications like voice cloning or translation-based TTS, as they map well to text-like representations such as LLM tokens. LLM-based [78, 103, 106, 105] and zero-shot TTS methods [17, 78, 87] often adopt discrete tokens as their acoustic features. However, discrete representation learning may result in information loss or lack the nuanced details that can be captured in continuous representations.
Table IV and III summarize the types of acoustic features of representative methods. Table II summarizes popular open-source speech quantization methods.
III Uncontrollable TTS
The development of Uncontrollable Text-To-Speech (UC-TTS) systems represents a significant shift from traditional, linguistics-based synthesis to modern, data-driven deep learning techniques. This shift highlights the integration of both local and global information to produce speech with human-like quality and naturalness. This survey explores UC-TTS evolution, emphasizing the role of local and global information in enhancing speech fidelity and expressiveness.
In the context of UC-TTS, ”uncontrollable” refers to the absence of explicit control mechanisms for speech features such as emotion, timbre, and speaker style. Despite this lack of explicit control, the goal is to achieve natural, fluid speech while minimizing issues like mispronunciations and omissions.
III-A Early Approaches: Statistical Models
Early Text-To-Speech (TTS) systems relied on statistical models such as Hidden Markov Models (HMMs) [64, 65] and early neural network-based parametric methods [111, 112]. These models operated at the frame level, using acoustic models and vocoders for text-to-speech conversion. Notable contributions from Tokuda et al. [173] employed HMMs for statistical parametric synthesis, focusing on local features like phonemes, accents, and prosody to improve speech naturalness.
While robust, these statistical methods were limited by their reliance on pre-segmented data, leading to oversimplified assumptions about speech dynamics. Local linguistic features were well-modeled, but the global phonetic context was often overlooked, resulting in speech that sounded monotone and lacked emotional depth, as noted by Zen et al. [174].
III-B Sequence-to-Sequence Models
The emergence of sequence-to-sequence models represents a significant breakthrough by removing the need for explicit linguistic features, thereby enabling the capture of the nuances and idiosyncrasies of human speech. Models such as Tacotron [74] and Tacotron 2  [175] utilize recurrent neural networks (RNNs) with attention mechanisms to effectively model the complex, nonlinear nature of speech sequences. These innovations allow for precise tuning of speech parameters, enhancing prosody and rhythm by modeling entire utterances rather than isolated phonetic units.
Building on these advancements, Deep Voice 3[176] introduces a fully convolutional sequence-to-sequence architecture that significantly accelerates training speed compared to RNN-based models. This approach achieves training times an order of magnitude faster, enabling scalability to handle large datasets. Additionally, the use of a position-augmented attention mechanism in Deep Voice 3 enhances the naturalness of synthesized speech, achieving competitive mean opinion scores, especially when paired with advanced neural vocoders like WaveNet. This development not only improves training efficiency but also enhances the scalability and naturalness of text-to-speech systems.
III-C Transformer-Based Models
Transformer-based architectures advanced the field by enabling computational parallelization and effectively capturing long-range dependencies. Models like Transformer TTS overcame RNN challenges, such as gradient vanishing, by using efficient training paradigms [177]. Self-attention mechanisms allowed simultaneous modeling of local phonetic details and global prosodic contexts, resulting in more sophisticated and human-like speech synthesis.
Although transformers improved contextual information incorporation, challenges remained in preserving local phonetic precision. To address these, techniques such as relative position encodings and localized attention were integrated [124].
III-D Advanced Architectures: Integrating Flow and Diffusion Models
Recent advancements have shifted towards integrating global information within end-to-end architectures to enhance speech naturalness and coherence. Flow-based models like Glow-TTS[133] and Flow-TTS[132] exemplify this by employing invertible transformations that maintain the balance between local precision and global coherence. These architectures enable the synthesis of high-fidelity speech by modeling complex dependencies across the entire utterance, thus improving the overall fluidity and naturalness of the generated speech.
Moreover, the introduction of diffusion models in TTS, such as WaveGrad 2[178], highlights the shift towards models that can iteratively refine speech output. These models use score matching and diffusion processes to generate speech directly from phoneme sequences, effectively capturing both local nuances and overarching global patterns. The iterative nature of these models allows for adjustments that enhance the quality of the synthesized audio, accommodating variations in speech without explicit control over specific attributes.
The integration of adversarial training and variational autoencoders (VAEs) further exemplifies the evolution towards incorporating global information. Systems like VITS[159] leverage these techniques to enhance expressiveness and naturalness by learning complex mappings between text and speech. This approach allows the model to manage variations in prosody and rhythm inherently derived from the textual input, aligning with the objectives of UC-TTS to produce diverse and natural speech outputs.
The evolution from HMMs to advanced architectures in UC-TTS exemplifies progress toward synthesizing speech that is both expressive and precise. The interplay of local and global information is crucial for enhancing speech quality and customizability. Future UC-TTS research aims to produce high-fidelity, customizable speech by harmonizing deep contextual insights with precise local adjustments, meeting diverse user needs and communication contexts.
IV Controllable TTS
In this section, we first review recent TTS work from the perspective of model architecture, followed by a detailed discussion of control modes in controllable TTS.
IV-A Model Architectures
Current model architectures can be broadly classified into two main categories: the first is the non-autoregressive (NAR) generative models, which are based on HMMs, neural networks, VAEs, diffusion models, flow matching, and other NAR techniques. The second category relies on autoregressive (AR) codec language models, which typically quantize speech into discrete tokens and use decoder-only models to autoregressively generate these tokens. We summarize the NAR-based and AR-based controllable TTS methods in Table III and Table IV, respectively.
IV-A1 Fully Non-Autoregressive (NAR) Architectures
HMM-based Approaches.  In the realm of Controllable Text-To-Speech (CTTS), advancements in Hidden Markov Model (HMM) architectures have significantly enhanced the manipulation of speech elements such as emotion and prosody. Yamagishi et al. [70] pioneered this field by introducing style-dependent and style-mixed modeling, which allowed precise emulation of human-like emotional nuances and versatile synthesis across various styles by incorporating style as a contextual variable. Building on this foundation, Qin et al.[179] developed the ”average emotion model,” which utilized MLLR-based adaptation to modulate emotions like happiness and sadness even with limited data, thus advancing the emotional intelligence of synthetic speech systems.
Furthering expressive variability, Nose et al.[119] integrated subjective style intensities and a multiple-regression global variance model into HMM frameworks, addressing over-smoothing and enabling nuanced emotional expressions. Lorenzo-Trueba et al.[72] expanded on these capabilities with CSMAPLR adaptation, introducing ”emotion transplantation” to transfer emotional states between speakers while preserving voice distinctiveness, enhancing personalized human-computer interaction. These innovations in HMM architectures have broadened the expressiveness and individuality in synthetic speech, augmenting technological interfaces and paving the way for future developments in adaptive, lifelike speech solutions.
Transformer-based Approaches.  Advancements in Controllable Text-to-Speech (TTS) technology highlight the integration of deep learning with audio processing, driven by Transformer-based architectures. Ren et al.[15] introduced FastSpeech, a feed-forward non-autoregressive Transformer model that significantly enhances TTS efficiency by reducing inference time and improving the stability issues found in autoregressive models like Tacotron 2. This model provides precise control over prosodic features through duration prediction, effectively tackling the one-to-many mapping challenge. FastSpeech 2[180] builds on this by integrating pitch and energy control, eliminating the need for the complex teacher-student distillation process, thus enhancing training efficiency and improving voice quality. Parallel Tacotron[84] further advances TTS by employing a variational autoencoder-based residual encoder, capturing intricate prosodic nuances. This approach, combined with iterative spectrogram loss, significantly enhances the naturalness and quality of synthesized speech. Additionally, FastPitch[77] incorporates direct pitch prediction into its architecture, enabling fully parallelized synthesis and precise pitch manipulation. This capability enhances expressiveness and retains the efficiency benefits established by FastSpeech. These innovations significantly contribute to the development of more interactive and natural AI-driven communication systems, underscoring the potential of integrating AI with human-centric disciplines to craft a future where technology and humanity coexist harmoniously.
VAE-based Approaches.  Recent advancements in Controllable Text-To-Speech (TTS) systems are largely driven by the integration of Variational Autoencoder (VAE) architectures, which enhance the flexibility and precision of speech modulation. Zhang et al.[18] pioneered the use of VAEs in end-to-end speech synthesis, creating disentangled latent representations that allow effective style control and transfer, especially in prosody and emotion management, outperforming the Global Style Token model in style transfer tasks. Building on this, Hsu et al.[130] developed a hierarchical generative model with a conditional VAE framework and a Gaussian mixture model, enabling precise control over complex speech attributes such as environment and style, thus improving expressive speech synthesis through refined noise and speaker characteristic management. Liu et al.[181] further advanced the field with the CLONE model, a single-stage TTS system that resolves the one-to-many mapping issue and enhances high-frequency information reconstruction. By employing a conditional VAE with normalizing flows and a dual path adversarial training mechanism with multi-band discriminators, CLONE achieves nuanced control over prosody and energy, demonstrating superior performance in both speech quality and prosody control compared to state-of-the-art models. These collective innovations highlight the adaptability of VAEs in managing complex speech generation tasks, marking significant progress toward more dynamic and versatile TTS technologies, with ongoing research promising even greater advancements.
Diffusion-based Approaches. The core concept of diffusion-based models is to generate target data by progressively removing noise. During the forward diffusion phase, noise is incrementally added to the original data to form a noise distribution. In the generation phase, a reverse denoising process is employed to gradually recover high-quality speech from the noise. Grad-StyleSpeech [182] introduces a hierarchical transformer encoder to create a representative noise prior distribution for speaker-adaptive settings using score-based diffusion models. NaturalSpeech 2 [86] uses a neural audio codec with residual vector quantizers to obtain quantized latent vectors, which are then generated using a diffusion model conditioned on text input. NaturalSpeech 3 [87] decomposes speech into distinct subspaces that represent different attributes and generates each subspace independently. DEX-TTS [183] improves DiT-based diffusion networks by applying overlapping patchify and convolution-frequency patch embedding strategies. E3 TTS [184] models the temporal structure of the waveform through the diffusion process, eliminating the need for any intermediate representations such as spectrogram features or alignment information.
Applying diffusion models to TTS requires a complex pipeline due to the need for precise temporal alignment between text and speech and the high fidelity required for audio data. This includes domain-specific modeling, such as phoneme and duration [86]. To address the issue of reduced naturalness caused by the addition of duration models, DiTTo-TTS [185] leverages the off-the-shelf pre-trained text and speech encoders without relying on speech domain-specific modeling by incorporating cross-attention mechanisms with the prediction of the total length of speech representations. Similarly, SimpleSpeech [186] proposes a speech codec model (SQ-Codec) based on scalar quantization and uses the sentence duration to control the generated speech length.
Flow-based Approaches. Flow-based methods leverage invertible flow transformations to learn mappings from target speech features to simple distributions, typically standard Gaussian distributions. Due to their invertibility, this mechanism can directly sample from the simple distribution and generate high-fidelity speech in the reverse direction. Audiobox [187] and P-flow [188] employ non-autoregressive flow-matching models for efficient and stable speech synthesis. VoiceBox [189] also employs a flow-matching to generate speech, effectively casting the TTS task into a speech infilling task. SpeechFlow [190] is trained on 60k hours of untranscribed speech with flow matching and mask conditions and can be fine-tuned with task-specific data to match or surpass existing expert models. This highlights the potential of generative models as foundation models for speech applications. HierSpeech++ [191] proposes a hierarchical variational inference method. FlashSpeech [192] is built on a latent consistency model and applies a novel adversarial consistency training approach that can train from scratch without the need for a pre-trained diffusion model as the teacher, achieving speech generation in one or two steps.
Recently, E2 TTS [193] converts text input into a character sequence with filler tokens and trains a mel spectrogram generator based on audio infilling task, achieving human-level naturalness. Inspired by E2 TTS, F5-TTS [194] refines the text representation with ConvNext v2 [195], facilitating easier alignment with speech. E1 TTS [196] further distills a diffusion-based TTS model into a one-step generator with distribution matching distillation [197, 198], reducing the number of network evaluations in sampling from diffusion models. SimpleSpeech 2 [199] introduces a flow-based scalar transformer diffusion model. The work also provides a theoretical analysis, showing that the inclusion of a small number of noisy labels in a large-scale dataset is equivalent to introducing classifier-free guidance during model optimization.
Other NAR Approaches. Other works leverage GAN-based or Masked Generative-based methods for TTS generation. StyleTTS 2 [89] employs large pre-trained speech language models (SLMs) such as Wav2Vec 2.0 [165], HuBERT [166], and WavLM [200] as discriminators, in combination with a novel differentiable duration modeling approach. This setup uses SLM representations to enhance the naturalness of the synthesized speech. MaskGCT [78] proposes masked generative transformers without requiring text-speech alignment supervision and phone-level duration. The model employs a two-stage system, both trained using a mask-and-predict learning paradigm.
IV-A2 Autoregressive (AR) Architectures
RNN-based Approaches. Controllable Text-To-Speech (TTS) technology has seen significant advancements through innovations in neural network architectures, facilitating speech that is both natural-sounding and adaptable in terms of emotion, prosody, and pitch. A key breakthrough was the introduction of Tacotron[74], a sequence-to-sequence architecture that effectively integrates prosodic variations, laying the groundwork for precise control over speech attributes. Tacotron 2[175] further enhanced this capability by better managing prosodic variability, though it averaged these variations, indicating a need for more sophisticated control methods. To address these constraints, Wang et al.[19] introduced Global Style Tokens (GSTs), using an unsupervised approach to encapsulate diverse speech styles into fixed tokens, thus enabling versatile style transfer within the Tacotron framework. Skerry-Ryan et al.[201] further advanced this by incorporating prosodic embeddings, providing detailed control over timing and intonation, and significantly improving the replication of emotional expressions in synthetic speech.
Building on these innovations, emotion-controllable models developed by Li et al.[21] focus on calibrating emotional nuances using emotion embedding networks and style loss alignment, allowing detailed modulation of emotional strength. Hierarchical models like MsEmoTTS[83] refine this approach by segmenting synthesis into global, utterance-level, and local emotional strengths, offering enhanced emotional expressiveness and intuitive control. These advancements have expanded the scope to produce nuanced TTS outputs, enabling precise control over emotion, prosody, and pitch, with applications ranging from virtual assistants to interactive narratives. As researchers continue to explore the potential of neural networks in TTS, the technology promises even richer, more engaging digital experiences, moving towards speech synthesis that is indistinguishable from natural human interaction.
LLM-based Approaches. Inspired by the success of large language models (LLMs) in natural language processing (NLP), recent studies have explored leveraging in-context learning for zero-shot TTS generation.
VALL-E [85] is a pioneering work in this area, formulating TTS as a conditional language modeling problem. It utilizes EnCodec [202] to discretize waveforms into tokens as intermediate representations and employs a two-stage modeling pipeline: an autoregressive model first generates coarse audio tokens, followed by a non-autoregressive model that iteratively predicts additional codebook codes for refinement. This hierarchical modeling of semantic and acoustic tokens has set the foundation for many subsequent LLM-based TTS approaches [203, 204, 205, 206].
Building on VALL-E, various improvements have been proposed. VALL-E X [207] extends VALL-E to multilingual scenarios, supporting zero-shot cross-lingual speech synthesis and speech-to-speech translation. ELLA-V [208] introduces a sequence order rearrangement step, enhancing local alignment between phoneme and acoustic modalities. RALL-E [209] incorporates prosody tokens as chain-of-thought prompting [210] to stabilize the generation of speech tokens. VALL-E R [211] improves phoneme-to-acoustic alignment and adopts codec-merging to boost decoding efficiency and reduce computational overhead. VALL-E 2 [212] introduces repetition aware sampling and grouped code modeling for greater stability and faster inference. HALL-E [213] adopts a hierarchical post-training framework, effectively managing the trade-off between reducing frame rate and producing high-quality speech.
Beyond the foundational improvements introduced by VALL-E and its immediate extensions, further advancements have focused on enhancing speech alignment, quality, and robustness. SpearTTS [203] and Make-a-voice [204] use semantic tokens to bridge the gap between text and acoustic features. FireRedTTS [214] further optimizes the tokenizer architecture to enhance speech quality. CoFi-Speech [215] generates speech in a coarse-to-fine manner via a multi-scale speech coding and generation approach, producing natural and intelligible speech. CosyVoice [17] employs supervised semantic tokens to enhance content consistency and speaker similarity in zero-shot voice cloning. Similarly, BASE TTS [216] introduces discrete speech representations based on the WavLM [200] self-supervised model, focusing on phonemic and prosodic information. SeedTTS [217] also proposes a self-distillation method for speech decomposition and a reinforcement learning approach to enhance the robustness, speaker similarity and controllability of generated speech. Based on this framework, Bailing-TTS [218] enhances the alignment of text and speech tokens using a continual semi-supervised learning strategy, enabling high-quality synthesis of Chinese dialect speech.
Although models using discrete tokens as intermediate representations have achieved notable success in zero-shot TTS, they still face fidelity issues compared to the continuous representation like Mel spectrograms [219, 220]. MELLE [219] optimizes the training objectives and sampling strategy, marking the first exploration of using continuous-valued tokens instead of discrete-valued tokens within the paradigm of autoregressive speech synthesis models. Similar to MELLE, ARDiT [220] encodes audio as vector sequence in continuous space and autoregressively generates these sequences by a decoder-only transformer.
TABLE III: A summary of existing non-autoregressive controllable neural-based methods.
Method Zero-shot TTS Controlability Model Architectures Acoustic Feature Release Time
Pit. Ene. Spe. Pro. Tim. Emo. Env. Des. Acoustic Model Vocoder
FastSpeech[15] Transformer WaveGlow MelS 2019.05
DWAPI [221] DNN Straight MelS + F0 + Intensity 2020.04
FastSpeech 2[180] Transformer Parallel WaveGAN MelS 2020.06
FastPitch[77] Transformer WaveGlow MelS 2020.06
Parallel Tacotron[222] Transformer + CNN WaveRNN MelS 2020.10
StyleTagging-TTS [223] Transformer + CNN HiFi-GAN MelS 2021.04
SC-GlowTTS [224] Transformer + Conv HiFi-GAN MelS 2021.06
Meta-StyleSpeech [225] Transformer MelGAN MelS 2021.06
DelightfulTTS[226] Transformer + CNN HiFiNet MelS 2021.11
YourTTS [82] Transformer HiFi-GAN LinS 2021.12
DiffGAN-TTS[227] Diffusion + GAN HiFi-GAN MelS 2022.01
StyleTTS [88] CNN + RNN + GAN HiFi-GAN MelS 2022.05
GenerSpeech [90] Transformer + Flow-based HiFi-GAN MelS 2022.05
NaturalSpeech 2 [86] Diffusion Codec Decoder Token 2022.05
Cauliflow[228] BERT + Flow UP WaveNet MelS 2022.06
CLONE[181] Transformer + CNN WaveNet MelS + LinS 2022.07
PromptTTS [101] Transformer HiFi-GAN MelS 2022.11
Grad-StyleSpeech [182] Score-based Diffusion HiFi-GAN MelS 2022.11
PromptStyle [229] VITS HiFi-GAN MelS 2023.05
StyleTTS 2 [89] Diffusion + GAN HifiGAN / iSTFTNet MelS 2023.06
VoiceBox [189] Flow Matching Diffusion HiFi-GAN MelS 2023.06
MegaTTS 2 [230] Diffusion + GAN HiFi-GAN MelS 2023.07
PromptTTS 2 [102] Diffusion Codec Decoder Token 2023.09
VoiceLDM [231] Diffusion HiFi-GAN MelS 2023.09
DuIAN-E[232] CNN + RNN HiFi-GAN MelS 2023.09
PromptTTS++ [233] Transformer + Diffusion BigVGAN MelS 2023.09
SpeechFlow [190] Flow Matching Diffusion HiFi-GAN MelS 2023.10
P-Flow [188] Flow Matching HiFi-GAN MelS 2023.10
E3 TTS [184] Diffusion / Waveform→Unet 2023.11
HierSpeech++ [191] Hierarchical Conditional VAE BigVGAN MelS 2023.11
Audiobox [187] Flow Matching EnCodec MelS 2023.12
FlashSpeech [192] Latent Consistency Model EnCodec Token 2024.04
NaturalSpeech 3 [87] Diffusion EnCodec Token 2024.04
InstructTTS [105] Transformer + Diffusion HiFi-GAN Token 2024.05
ControlSpeech [106] Transformer + Diffusion FACodec Decoder Token 2024.06
AST-LDM [234] Diffusion HiFi-GAN MelS 2024.06
SimpleSpeech [186] Transformer Diffusion SQ Decoder Token 2024.06
DiTTo-TTS [185] DiT BigVGAN Token 2024.06
E2 TTS [193] Flow Matching Transformer BigVGAN MelS 2024.06
MobileSpeech [235] ConFormer Decoder Vocos Token 2024.06
DEX-TTS [183] Diffusion HiFi-GAN MelS 2024.06
ArtSpeech [236] RNN + CNN HiFI-GAN MelS+Energy+F0+TV 2024.07
CCSP [237] Diffusion Codec Decoder Token 2024.07
SimpleSpeech 2 [199] Flow-based Transformer Diffusion SQ Decoder Token 2024.08
E1 TTS [196] DiT BigVGAN Continuous Token 2024.09
VoiceGuider [238] Diffusion BigVGAN MelS 2024.09
StyleTTS-ZS [239] Diffusion + GAN HifiGAN / iSTFTNet Token 2024.09
NansyTTS [240] Transformer NANSY++ MelS 2024.09
NanoVoice [241] Diffusion BigVGAN MelS 2024.09
MS2KU-VTTS [242] Diffusion BigvGAN MelS 2024.10
MaskGCT [78] Masked Generative Transformers DAC + Vocos Token 2024.10
Abbreviations: Pit(ch), Ene(rgy), Spe(ed), Pro(sody), Tim(bre), Emo(tion), Env(ironment), Des(cription). Timbre involves gender and age. MelS and LinS represent Mel Spectrogram and Linear Spectrogram respectively. TV represents the vocal tract variables proposed in  [236].
TABLE IV: A summary of existing autoregressive controllable neural-based methods.
Method Zero-shot TTS Controlability Model Architectures Acoustic Feature Release Time
Pit. Ene. Spe. Pro. Tim. Emo. Env. Des. Acoustic Model Vocoder
Prosody-Tacotron[201] RNN WaveNet MelS 2018.03
GST-Tacotron[243] CNN + RNN Griffin-Lim LinS 2018.03
GMVAE-Tacotron[130] CNN + RNN WaveRNN MelS 2018.12
VAE-Tacotron[18] CNN + RNN WaveNet MelS 2019.02
DurIAN[244] CNN + RNN MB-WaveRNN MelS 2019.09
Flowtron[245] CNN + RNN WaveGlow MelS 2020.07
MsEmoTTS[83] CNN + RNN WaveRNN MelS 2022.01
VALL-E [85] LLM EnCodec Token 2023.01
SpearTTS [203] LLM SoundStream Token 2023.02
VALL-E X [207] LLM EnCodec Token 2023.03
Make-a-voice [204] LLM BigVGAN Token 2023.05
TorToise [246] Transformer + DDPM Univnet MelS 2023.05
MegaTTS [91] LLM + GAN HiFi-GAN MelS 2023.06
SC VALL-E [205] LLM EnCodec Token 2023.07
Salle [247] LLM Codec Decoder Token 2023.08
UniAudio [248] LLM EnCodec Token 2023.10
ELLA-V [208] LLM EnCodec Token 2024.01
BaseTTS [216] LLM UnivNet Token 2024.02
ClaM-TTS [249] LLM BigVGAN MelS+Token 2024.04
RALL-E [209] LLM SoundStream Token 2024.05
ARDiT [220] Decoder-only Diffusion Transformer BigVGAN MelS 2024.06
VALL-E R [211] LLM Vocos Token 2024.06
VALL-E 2 [212] LLM Vocos Token 2024.06
Seed-TTS [217] LLM + Diffusion Transformer / Token 2024.06
VoiceCraft [93] LLM HiFi-GAN Token 2024.06
XTTS [250] LLM + GAN HiFi-GAN MelS+Token 2024.06
CosyVoice [17] LLM + Conditional Flow Matching HiFi-GAN Token 2024.07
MELLE [219] LLM HiFi-GAN MelS 2024.07
Bailing TTS [218] LLM + Diffusion Transformer / Token 2024.08
VoxInstruct [103] LLM Vocos Token 2024.08
Emo-DPO [31] LLM HiFi-GAN Token 2024.09
FireRedTTS [214] LLM + Conditional Flow Matching BigVGAN-v2 Token 2024.09
CoFi-Speech [215] LLM BigVGAN Token 2024.09
Takin [206] LLM HiFi-Codec Token 2024.09
HALL-E [213] LLM EnCodec Token 2024.10
Abbreviations: Pit(ch), Ene(rgy), Spe(ed), Pro(sody), Tim(bre), Emo(tion), Env(ironment), Des(cription). Timbre involves gender and age. MelS and LinS represent Mel Spectrogram and Linear Spectrogram respectively.
IV-B Control Strategies
The control strategies in existing controllable TTS can be broadly classified into four categories: style tagging using discrete labels, speech reference prompt for customizing a new speaker’s voice with just a few seconds of voice input, controlling speech style using natural language descriptions, and the instruction-guided mode. We illustrate taxonomies of controllable TTS from the perspective of control strategies in Fig. 3.
[…]
Figure 3: A taxonomy of controllable TTS from the perspective of control strategies.
IV-B1 Style Tagging
This paradigm typically employs target control attributes, primarily emotion-related controls, as categorical label inputs to enable controllable speech synthesis. StyleTagging-TTS [223] utilizes a short phrase or word to represent the style of an utterance and learns the relationship between linguistic embedding and style embedding space by a pre-trained language model.
However, these methods are limited in expressive diversity, as they can only model a small set of pre-defined styles.
IV-B2 Reference Speech Prompt
This paradigm aims to customize a new speaker’s voice with just a few seconds of voice prompt. The architecture can be abstracted into two main components: a speaker encoder that processes the reference speech and outputs a speaker embedding, and a conditional TTS decoder that takes both text and speaker embedding as input to generate speech that matches the style of the reference prompt. MetaStyleSpeech [225] and StyleTTS [88] use adaptive normalization as a style conditioning method, enabling robust zero-shot performance. GenerSpeech [90] introduces a multi-level style adapter to improve zero-shot style transfer for out-of-domain custom voices. SC VALL-E [205] facilitates control over synthesized speech’s emotions, speaking styles, and various acoustic features by incorporating style tokens and scale factors. ArtSpeech [236] revisits the sound production system by integrating articulatory representations into the TTS framework, improving the physical interpretability of articulation movements.
To enhance the learning of contextual information and address the challenge of limited voice data from the target speaker, CCSP [237] proposes a contrastive context-speech pretraining framework that learns cross-modal representations combining both contextual text and speech expressions. DEX-TTS [183] separates styles into time-invariant and time-variant components, enabling the extraction of diverse styles from expressive reference speech. StyleTTS-ZS [239] leverages distilled time-varying style diffusion to capture diverse speaker identities and prosodies.
Some works also decouple timbre and style information from the reference speech, allowing more flexible control over the speaking style [87, 230, 106]. MegaTTS 2 [230] introduces an acoustic autoencoder that separately encodes prosody and timbre into the latent space, enabling the transfer of various speaking styles to the desired timbre. ControlSpeech [106] uses bidirectional attention and mask-based parallel decoding to capture codec representations in a discrete decoupling codec space, allowing independent control of timbre, style, and content in a zero-shot manner.
IV-B3 Natural Language Descriptions
Recent studies explore controlling speech style using natural language descriptions that include attributes such as pitch, gender, and emotion, making the process more user-friendly and interpretable. In this paradigm, several speech datasets with natural language descriptions [101, 247, 106] and associated prompt generation pipelines [247, 102, 252] have been proposed. Detailed information about these datasets will be discussed in Section V. PromptTTS [101] uses manually annotated text prompts to describe five speech attributes, including gender, pitch, speaking speed, volume, and emotion. InstructTTS [105] introduces a three-stage training procedure to capture semantic information from natural language style prompts and adds further annotation to the NLSpeech dataset’s speech styles. PromptStyle [229] constructs a shared space for stylistic and semantic representations through a two-stage training process. TextrolSpeech [247] proposes an efficient prompt programming methodology and a multi-stage discrete style token-guided control framework, demonstrating strong in-context capabilities. NansyTTS [240] combines a TTS trained on the target language with a description control model trained on another language, which shares the same timbre and style representations to enable cross-lingual controllability.
Considering not all details about voice variability can be described in the text prompt, PromptTTS++ [233] and PromptSpeaker [251] tries to construct text prompts with more details. PromptTTS 2 [102] designs a variation network to capture voice variability not conveyed by text prompts. ControlSpeech [106] proposes the Style Mixture Semantic Density (SMSD) module, incorporating a noise perturbation mechanism to tackle the many-to-many problem in style control and enhance style diversity.
Other works also focus on improving controllability in additional aspects, such as the surrounding environment. Audiobox [187] introduces both description-based and example-based prompting, integrating speech and sound generation paradigms to independently control transcript, vocal, and other audio styles during speech generation. VoiceLDM [231] and AST-LDM [234] extend AudioLDM [253] to incorporate environmental context in TTS generation by adding a content prompt as a conditional input. Building on VoiceLDM, MS2KU-VTTS [242] further expands the dimensions of environmental perception, enhancing the generation of immersive spatial speech.
IV-B4 Instruction-Guided Control
The natural language description-based TTS methods discussed above require splitting inputs into content and description prompts, which limits fine-grained control over speech and does not align with other AIGC models. VoxInstruct [103] proposes a new paradigm that extends traditional text-to-speech tasks into a general human instruction-to-speech task. Here, human instructions are freely written in natural language, encompassing both the spoken content and descriptive information about the speech. To enable automatic extraction of the synthesized speech content from raw text instructions, VoxInstruct uses speech semantic tokens as an intermediate representation, bridging the gap in current research by allowing the simultaneous use of both text description prompts and speech prompts for speech generation.
V Datasets and Evaluation
TABLE V: A summary of open-source datasets for controllable TTS.
Dataset Hours #Speakers Labels Lang Release Time
Pit. Ene. Spe. Age Gen. Emo. Emp. Acc. Top. Des. Env. Dia.
Taskmaster-1 [254] / / en 2019.09
Libri-light [255] 60,000 9,722 en 2019.12
AISHELL-3 [256] 85 218 zh 2020.10
ESD [257] 29 10 en,zh 2021.05
GigaSpeech [258] 10,000 / en 2021.06
WenetSpeech [259] 10,000 / zh 2021.07
PromptSpeech [101] / / en 2022.11
DailyTalk [260] 20 2 en 2023.05
TextrolSpeech [247] 330 1,324 en 2023.08
VoiceLDM [231] / / en 2023.09
VccmDataset [106] 330 1,324 en 2024.06
MSceneSpeech [261] 13 13 zh 2024.07
SpeechCraft [262] 2,391 3,200 en,zh 2024.08
Abbreviations: Pit(ch), Ene(rgy)=volume=loudness, Spe(ed)=duration, Gen(der), Emo(tion), Emp(hasis), Acc(ent), Dia(logue), Env(ironment), Des(cription).
TABLE VI: Common objective and subjective evaluation metrics
Metric Type Eval Target GT Required
MCD [263] Objective Acoustic similarity
PESQ [264] Objective Perceptual quality
WER [265] Objective Intelligibility
MOS [266] Subjective Preference
CMOS [267] Subjective Preference
GT: Ground truth.
V-A Datasets
Achieving fully controllable TTS requires large-scale datasets with rich diversity and fine-grained annotations. In this subsection, we categorize speech datasets into four types according to the labels they provide, i.e., attribute tags such as age and gender, description, environment, and dialogues.
Tag-based speech datasets. Tag-based datasets [255, 256, 257, 258, 259, 260, 261] are specialized collections of speech data that include metadata tags, such as pitch, emotion, and energy, age, gender, and so on, to enhance the expressiveness and control of TTS systems. These datasets not only provide audio-text pairs but also offer additional labels that guide the model in generating diverse and context-aware speech outputs.
Description-based speech datasets. Description-based TTS datasets [101, 247, 231, 106, 262] contain speech data paired with detailed textual descriptions of the speech attributes or characteristics, such as emotion, speed, intonation, and style. These datasets enable training TTS models that interpret descriptive prompts to generate expressive and context-aware speech.
Speech environment datasets. Speech environment datasets [231, 268] are collections of speech data annotated with environmental labels, such as park, stadium, office, or street, capturing the acoustic characteristics of various real-world settings. These datasets are crucial for training models to handle diverse acoustic scenarios, improving their robustness and adaptability.
Dialogue speech datasets. Dialogue speech datasets [260, 254] capture conversational speech between two or more people, focusing on the natural flow, turn-taking, and context of human dialogue. These datasets are crucial for training systems in applications like chatbots, virtual assistants, and interactive storytelling.
We summarize open-source speech datasets for controllable TTS in Table V.
V-B Evaluation
The performance of controllable TTS often requires objective and subjective evaluation. We introduce common evaluation metrics in this subsection.
Objective evaluation metrics. Objective metrics offer automated and reproducible evaluations. Mel Cepstral Distortion (MCD) [263] measures the spectral distance between synthesized and reference speech, reflecting how closely the generated audio matches the target in terms of acoustic features. For intelligibility, the Word Error Rate (WER) [265] is used, comparing transcriptions of synthesized speech to the input text via automated speech recognition. Perceptual Evaluation of Speech Quality (PESQ) [264] is another objective metric designed to evaluate speech quality by comparing degraded audio with a clean reference. It is widely used in telecommunications and speech synthesis, PESQ models human auditory perception, producing a score (typically 1–4.5) that reflects intelligibility and distortion under various conditions, including noise or compression.
Subjective evaluation metrics. The Mean Opinion Score (MOS) [266] is the most commonly used subjective metric. In MOS evaluations, listeners rate the naturalness of synthesized speech on a scale (e.g., 1 to 5), where higher scores indicate better quality. MOS captures human perception effectively but requires substantial resources for large-scale evaluations. Comparison Mean Opinion Score (CMOS) [267] further evaluates relative quality differences between two TTS audio samples. Participants listen to paired samples and rate their preference on a scale (e.g., -3 to +3, where negative values favor the first sample). CMOS is used to measure subtle improvements in TTS systems, complementing absolute MOS ratings.
Table VI summarizes common evaluation metrics for TTS.
VI Challenges and Future Directions
In this section, we elaborate on current challenges for fully controllable TTS and discuss promising future directions.
VI-A Challenges
Controllable TTS aims to synthesize speech while allowing precise control over speech characteristics such as pitch, duration, energy, prosody, speaking style, and emotion. While significant progress has been made, achieving truly controllable TTS remains a complex task due to the multifaceted nature of human speech and the technical challenges in modeling and synthesizing it. In this section, we delve into the primary challenges and analyze their underlying reasons.
Controllable granularity. A critical challenge in controllable TTS is determining what aspects of speech should be controlled and how to control speech characteristics at a specific granularity. Different applications require varying levels of control granularity. For instance, audiobook narration may need sentence-level control of emotion, while conversational AI like ChatGPT may require word or phoneme-level control over prosody. Moreover, the emotion, prosody, and other characteristics of human speech are often intricately intertwined and can manifest across varying levels of granularity. Additionally, achieving fine-grained control requires high-resolution annotations and sophisticated models capable of handling subtle variations without compromising synthesis quality.
Although some LLM-based TTS methods such as VoxInstruct [103] can control various aspects of speech through attribute descriptions, determining the appropriate level of granularity for control and devising methods to achieve precise control at a specific granularity or to enable multiscale and fine-grained control remains a significant challenge.
Feature extraction and representation. Achieving fully controllable TTS needs good feature disentanglement. Accurately extracting meaningful and disentangled speech features like pitch contours, energy patterns, emotion variation, and prosodic elements from training data is difficult. The reason is that speech features are interdependent and context-sensitive, making it hard to isolate specific attributes for control. For example, altering pitch often affects prosody, emotion, and naturalness to some extent. To tackle this, several methods [269, 270, 271] utilize pre-trained models for different speech recognition tasks (e.g., pitch, energy, and duration prediction, gender classification, age estimation, and speaker verification) to supervise feature extraction. For example, NaturalSpeech3 [14] factorizes speech into separate feature subspaces to capture different speech attributes.
However, these methods are limited to coarse or high-level feature disentanglement, leaving a significant gap in fully disentangled control. On the other hand, selecting suitable representations (e.g., continuous variables like mel-spectrograms or latent embeddings like tokens) for controllable attributes is non-trivial because representations must be both interpretable for humans and expressive enough for TTS models. For example, transformer-based models are good at processing discrete tokens, while GAN and Diffusion-based models excel in modeling continuous representations.
Scarcity of datasets. High-quality, diverse, and appropriately annotated datasets are essential for training controllable TTS systems. However, such datasets are scarce and difficult to construct. To achieve controllable TTS, training data must encompass a wide range of styles, emotions, accents, and prosodic variations to enable versatile control because limited diversity in datasets can restrict the model’s ability to generalize across unseen styles or emotions. Although there are some TTS datasets, such as LibriTTS [272], Gigaspeech [258], and TextrolSpeech [247], their diversity is still not enough for fully controllable TTS due to the lack of corpus of diverse content such as comedies, thrillers, cartoons, etc. Constructing large-scale datasets with rich diversity is also expensive and time-consuming.
Another obstacle is that creating datasets with fine-grained, attribute-specific annotations is labor-intensive and costly. Besides, manual annotation of speech attributes requires expert knowledge and is prone to inconsistencies and errors, particularly for subjective qualities like emotion. Currently, most datasets provide only coarse labels, such as gender, age, or a limited range of emotions. While some datasets, such as SpeechCraft [262], include natural language descriptions of speech attributes, no existing dataset offers fine-grained variations and annotations within the speech of the same speakers. Available datasets for controllable TTS are summarized in Table V.
Generalization ability. The ability of a TTS system to generalize effectively is crucial for producing natural, high-quality speech across a wide range of conditions, such as unseen speakers, languages, or topics. However, achieving robust generalization remains a significant challenge for modern TTS methods due to various factors. Zero-shot controllable TTS [78, 92] aims to synthesize speech for unseen speakers with various speech customization such as emotion using minimal reference audio, which can offer flexibility for personalized voice generation. However, it faces significant challenges, including capturing unique speaker characteristics from limited data, accurately reproducing prosody and style, and disentangling speaker identity from other audio attributes like emotion or noise.
Multilingual generalization [250, 273] in TTS refers to the ability to synthesize natural and intelligible speech across multiple languages, including those not seen during training. This capability is essential for applications like cross-lingual communication, multilingual virtual assistants, and speech synthesis for low-resource languages [274]. Multilingual generalization still faces many challenges such as linguistic diversity and mismatch and the scarcity of data. Cross-lingual speaker generalization is another hurdle, as preserving speaker identity across languages can lead to artifacts.
Domain adaptation [275] in TTS refers to tailoring a pre-trained TTS model to generate speech for a specific domain or context, such as medical terminology and conversational speech. One challenge is that many specialized domains lack sufficient high-quality annotated data for fine-tuning. Besides, adapting prosody, intonation, and speaking style to match domain-specific requirements such as comic dialogue is complex. Failing to capture domain-specific nuances can make speech sound unnatural or inconsistent with the target context.
Efficiency. Efficiency in controllable TTS systems is a critical requirement for practical applications, as these models aim to offer fine-grained control over various speech attributes such as prosody, emotion, style, and speaker identity. However, achieving such control often comes at the cost of increased computational complexity, larger model sizes, and longer inference times, creating significant challenges.
High latency is a major issue, as existing controllable TTS models [78, 101, 102, 103] often necessitate autoregressive processes to synthesize speech. The inference time of these models ranges from several to tens of seconds. This can be especially problematic for real-time applications like live broadcasting or interactive systems. Additionally, the challenge of balancing granularity and efficiency arises, as finer controls demand higher-resolution data and more precise models, leading to increased resource requirements and inefficient training and inference.
Another major obstacle lies in the trade-off between model complexity and performance. State-of-the-art controllable TTS systems often rely on large neural networks such as LLMs with billions of parameters, which provide superior naturalness and expressiveness but demand significant computational resources. Simplifying these architectures can lead to quality degradation, including artifacts, unnatural prosody, or limited expressiveness. Therefore, designing light-weight controllable TTS models is significantly tricky.
[…]
Figure 4: A summary of current major challenges for fully controllable TTS.
VI-B Future Directions
In this survey, we conduct a comprehensive investigation and analysis of existing TTS methods, particularly on controllable TTS technologies.  While these methods show great potential in real-world applications, there are still some limitations that need to be addressed.  Based on our observations, we outline several promising future directions as follows:
Fine-grained speech synthesis by natural language description. Using natural language description to synthesize human speech with fine-grained control over various audio attributes is currently underexplored. Most of the existing works can only control a fixed number of attributes of the synthesized speech. Although a few works show great control of emotion, timbres, pitch, gender, and styles, e.g., VoxInstruct [103] and CosyVoice [17], they can frequently synthesize unwanted speech clips. Users need to synthesize multiple times to get satisfactory speech.
Fine-grained speech editing by natural language description. Speech or audio editing has been studied for a long time. However, existing methods usually train conditional models and adjust a fixed number of conditional inputs to modify the attributes of synthesized speech, thus lacking fine-grained manipulations [95, 94]. Therefore, how to learn disentangled speech representations for speech attributes while supporting editing by using natural language description is worthy of investigation.
Expressive multi-modal speech synthesis. Synthesizing speech from multi-modal data such as texts, images, and videos is an appealing research topic due to its various applications in the industry such as storytelling, filming, and gaming. Although there are several related works on this task [6, 24, 276, 277], few of them can fully extract useful information from multi-modal data. Particularly, synthesizing engaging speech and expressive voiceover for complex visual content sees great opportunities in the future.
Natural and emotional conversational TTS. Speech conversational TTS have come out for several decades but remained as cascaded systems for a long time and cannot generate natural and emotional speech. These systems are not context-aware, making the synthesized speech sound robotic. With the advent of LLMs, existing TTS technologies were directly introduced by simply synthesizing speech from the text generated by LLMs [33]. However, context-aware conversational TTS with rich emotion and good naturalness has not been well studied.
Zero-shot long speech synthesis with emotion consistency. Zero-shot TTS emerged in recent years to achieve voice cloning and speech style imitation without fine-tuning, making them more practical in real scenarios [78, 194, 17]. However, synthesizing long speech with rich emotion and style variation in a zero-shot setting remains challenging due to the lack of rich speech information in short reference audio clips. Addressing this issue will make a big step towards fully controllable zero-shot TTS.
Efficient TTS by natural language description. Synthesizing speech with natural language description usually involves training large language encoders and bridge nets between the two modalities which can bring about much more computation overhead compared to previous TTS methods. The inference time is also relatively slow, e.g., existing methods usually take tens of seconds to synthesize a short speech audio clip of less than 10 seconds [17, 104]. Therefore, efficient text and speech modeling and interaction is critical for natural language description-based TTS systems.
VII Discussion
VII-A Impacts of Controllable TTS
Applications. Controllable TTS systems allow fine-grained manipulation of speech attributes such as pitch, emotion, speaking rate, and style, enabling a wide range of applications across industries. One major application is in virtual assistants and customer support systems, where controllable TTS ensures tailored and context-aware responses. For instance, a virtual assistant can speak in a calm tone for technical support but switch to an enthusiastic tone when presenting promotional offers, enhancing user experience.
In the entertainment industry, controllable TTS is invaluable for creating dynamic voiceovers, audiobooks, and gaming characters. It enables precise adjustments in tone and delivery, allowing audiobooks to reflect character emotions and gaming characters to exhibit personality traits that align with their roles. Similarly, in education, TTS systems can adapt speaking styles to suit different learners, such as adopting a slow, clear tone for language learning or an engaging, storytelling style for children’s content.
Controllable TTS is also transformative in assistive technologies, where it can generate speech that reflects the user’s intended emotion or personality. This is particularly impactful for individuals with speech impairments, enabling more expressive and natural communication.
In content localization, controllable TTS systems adjust speech characteristics to suit regional and cultural preferences, ensuring a natural fit for global audiences. Additionally, it finds applications in human-computer interaction research, enabling adaptive dialogue systems that modify speech dynamically based on user mood or environment. By offering this flexibility, controllable TTS systems enhance accessibility, personalization, and engagement across various domains.
Deepfakes. A deepfake is a type of synthetic media in which a person in an existing image, video, or audio recording is replaced with someone else’s likeness or voice. This technology uses deep learning, particularly GANs [278], to create highly realistic, but fabricated, content. While deepfakes are most commonly associated with video manipulation, such as face swapping [279], they can also apply to audio, enabling the creation of synthetic speech that mimics a specific person’s voice, which is well known as voice cloning.
Voice cloning, especially few-shot [280] and zero-shot TTS [85, 78], poses a significant threat to systems that rely on voice authentication, such as banking, customer service, and other identity verification processes. With a convincing clone of someone’s voice, attackers could potentially impersonate individuals to gain unauthorized access to sensitive information or accounts.
To address these concerns, it’s essential to establish robust security protocols, consent-based regulations, and public awareness around voice cloning. Furthermore, advancements in detecting voice clones are equally important to help distinguish genuine voices from synthesized ones, protecting both individuals and organizations from potential misuse.
VII-B Limitation of this Survey
Although we conduct a comprehensive survey of controllable TTS in this paper, there are some limitations we want to address in the future: 1) A unified benchmark method will be provided to evaluate controllable TTS methods. 2) Detailed analysis and control strategies of each specific speech attribute will be provided in an updated version of this paper. 3) The methodologies for feature disentanglement are crucial for controllable TTS but are not adequately discussed.
VIII Conclusion
In this survey paper, we first elaborate on the general pipeline for controllable TTS, followed by a glimpse of uncontrollable TTS methods. Then we comprehensively review existing controllable methods from the perspectives of model architectures and control strategies. Popular datasets and commonly used evaluation metrics for controllable TTS are also summarized in this paper. Besides, the current challenges are deeply analyzed and the promising future directions are also pointed out. To the best of our knowledge, this is the first comprehensive survey for controllable TTS.
Writing a comprehensive survey is a challenging task and an iterative process. Hence, we will regularly update this survey to offer researchers and practitioners a continuously evolving resource for understanding controllable speech synthesis.

### v2

@@
 This includes fine-grained control over various attributes of synthesized speech such as emotion, prosody, timbre, and duration.
-Besides, advancements in deep learning, such as diffusion and large language models, have significantly enhanced controllable TTS over the past several years.
-In this paper, we conduct a comprehensive survey of controllable TTS, covering approaches ranging from basic control techniques to methods utilizing natural language prompts, aiming to provide a clear understanding of the current state of research.
+In addition, advancements in deep learning, such as diffusion and large language models, have significantly enhanced controllable TTS over the past several years.
+In this work, we conduct a comprehensive survey of controllable TTS, covering approaches ranging from basic control techniques to methods utilizing natural language prompts, aiming to provide a clear understanding of the current state of research.
 We examine the general controllable TTS pipeline, challenges, model architectures, and control strategies, offering a comprehensive and clear taxonomy of existing methods.
 Additionally, we provide a detailed summary of datasets and evaluation metrics and shed some light on the applications and future directions of controllable TTS.
-To the best of our knowledge, this survey paper provides the first comprehensive review of emerging controllable TTS methods, which can serve as a beneficial resource for both academic researchers and industry practitioners.
+To the best of our knowledge, this survey paper provides the first comprehensive review of emerging controllable TTS methods, which can serve as a beneficial resource for both academic researchers and industrial practitioners.
 Index Terms:
@@
 Meanwhile, the ability to achieve fine-grained control over synthesized speech attributes, such as emotion, prosody, timbre, and duration, has become a hot research topic in both academia and industry, driven by its vast potential for diverse applications.
-Deep learning [12] has made great progress in the past decade due to exponentially growing computational resources like GPUs [13], leading to the explosion of numerous great works on TTS [14, 15, 16, 17].
-These methods can synthesize human speech with better quality [14] and can achieve fine-grained control of the generated voice [18, 19, 20, 21, 22].
-Besides, some recent works synthesize speech given multi-modal input, such as face images [23, 24], cartoons [7], and videos [25].
-Moreover, with the fast development of open-source LLMs [11, 26, 27, 28, 29], some researchers propose to synthesize fine-grained controllable speech with natural language description [30, 31, 32], coining a new way to generate custom speech voices.
+Deep learning [12] has made great progress in the past decade due to exponentially growing computational resources like GPUs [13], leading to the explosion of numerous exciting works on TTS [14, 15, 16, 17].
+These methods can synthesize human speech with improved quality [14] and can achieve fine-grained control of the generated voice [18, 19, 20, 21, 22].
+In addition, some recent works synthesize speech given multi-modal input, such as face images [23, 24], cartoons [7], and videos [25].
+Moreover, with the fast development of open-source LLMs [11, 26, 27, 28, 29], some researchers propose to synthesize fine-grained controllable speech with natural language description [30, 31, 32], offering a new way to generate custom speech voices.
 Meanwhile, powering LLMs with speech synthesis has also been a hot topic in the last few years [33, 34, 35].
-In recent years, a wide range of TTS methods has emerged, making it essential for researchers to gain a comprehensive understanding of current research trends, particularly in controllable TTS, to identify promising future directions in this rapidly evolving field.
+In recent years, a wide range of TTS methods has emerged, making it essential for researchers to gain a comprehensive understanding of current research trends, particularly in controllable TTS, and to identify promising future directions in this rapidly evolving field.
 Consequently, there is a pressing need for an up-to-date survey of TTS techniques.
-While several existing surveys address parametric-based approaches [36, 37, 38, 39, 40, 41] and deep learning-based TTS [42, 43, 44, 45, 46, 47, 48], they largely overlook the controllability of TTS.
-Additionally, these surveys do not cover the advancements in recent years, such as natural language description-based TTS methods.
+While several existing surveys address parametric approaches [36, 37, 38, 39, 40, 41] and deep learning-based approaches [42, 43, 44, 45, 46, 47, 48], they largely overlook the controllability of TTS.
+Additionally, these surveys do not cover recent advancements, such as natural language description-based TTS methods.
 This paper provides a comprehensive and in-depth survey of existing and emerging TTS technologies, with a particular focus on controllable TTS methods.
@@
 Finally, we introduce the taxonomy and organization of this paper.
+We have posted a version of our paper on arXiv.org (https://arxiv.org/abs/2412.06602).
 Figure 1: A summary of representative controllable TTS methods in recent years and their model architectures, feature representations, and control abilities.
-There are additional network structures, such as VAE-based models, that are not included in this figure.
+Additional network structures, such as VAE and flow-based models, are not included in this figure.
 For more details, refer to Tables IV and III.
@@
 [49] and King et al.
-[40] explored rule-based, concatenative, and HMM-based techniques.
-Later, the advent of deep learning catalyzed the emergence of numerous neural-based TTS methods.
+[40] explored rule-based, concatenative, and Hidden Markov Models (HMM)-based techniques.
+Later, the advent of deep learning catalyzed the emergence of numerous neural model-based TTS methods.
 Therefore, Ning et al.
 [43] and Tan et al.
-[42] have conducted extensive surveys on neural-based acoustic models and vocoders, while Zhang et al.
+[42] have conducted extensive surveys on neural acoustic models and vocoders, while Zhang et al.
 [50] presented the first review of diffusion model-based TTS techniques.
@@
 To address this gap, we present the first comprehensive survey of TTS methods through the lens of controllability, providing an in-depth analysis of model architectures and strategies for controlling synthesized speech.
-Close to current demand.
-With the rapid development of hardware (i.e., GPUs) and artificial intelligence techniques (i.e., transformers, LLMs, diffusion models) in the last few years, the demand for controllable TTS is becoming increasingly urgent due to its broad applications in industries such as filmmaking, gaming, robots, and personal assistants.
+Close to Current Demands.
+With the rapid development of hardware (i.e., GPUs and TPUs) and artificial intelligence (AI) techniques (i.e., transformers, LLMs, diffusion models) in the last few years, the demand for controllable TTS is becoming increasingly urgent due to its broad applications in industries such as filmmaking, gaming, robotics, and personal assistants.
 Despite this growing need, existing surveys pay little attention to control methods in TTS technologies.
@@
 Based on these analyses, we identify promising directions for future research on TTS technologies.
-Table I summarizes representative surveys and this paper in terms of main focus and publication year.
+Table I summarizes the existing surveys and our survey in terms of main focus and publication year.
 TABLE I: Comparison with representative TTS surveys.
@@
 King et al. [40] Parametric TTS and performance measurement 2014
-Tan et al. [42] Neural-based, efficient, and expressive TTS 2021
+Tan et al. [42] Neural, efficient, and expressive TTS 2021
 Zhang et al. [50] Diffusion-based TTS and speech enhancement 2023
-Ours Controllable TTS and evaluation 2024
+Ours Controllable TTS, datasets, metrics, and challenges 2025
 I-B The History of Controllable TTS
 Controllable TTS aims to control various aspects of synthesized speech, such as pitch, energy, speed/duration, prosody, timbre, emotion, gender, or high-level styles.
-This subsection briefly reviews the history of controllable TTS ranging from early approaches to the state-of-arts in recent years.
+This subsection briefly reviews the history of controllable TTS, ranging from early approaches to the state-of-the-art (SOTA) in recent years.
 Early Approaches.
 Before the prevalence of deep neural networks (DNNs), controllable TTS technologies were built primarily on rule-based, concatenative, and statistical methods.
-These approaches enable some degree of customization and control, though they were constrained by the limitations of the underlying models and available computational resources.
+These approaches enable some degree of customization and control, however, they were constrained by the limitations of the underlying models and available computational resources.
 1) Rule-based TTS systems [51, 52, 53, 54], such as formant synthesis, were among the earliest methods for speech generation.
@@
 Some HMM-based systems also supported speaker adaptation [66, 67] and voice conversion [68, 69], enabling voice cloning to some extent.
-Besides, emotion can also be limitedly controlled by some of these methods [70, 71, 60, 72].
+However, emotion can be limitedly controlled by some of these methods [70, 71, 60, 72].
 In addition, they required less storage compared to concatenative TTS and allowed smoother transitions between speech units.
-Neural-based synthesis.
-Neural-based TTS technologies emerged with the advent of deep learning, significantly advancing the field by enabling more flexible, natural, and expressive speech synthesis.
-Unlike traditional methods, neural-based TTS leverages DNNs to model complex relationships between input text and speech, facilitating nuanced control over various speech characteristics.
-Early neural TTS systems, such as WaveNet [73] and Tacotron [74] laid the groundwork for controllability.
+Neural Synthesis.
+Neural model-based TTS technologies emerged with the advent of deep learning, significantly advancing the field by enabling more flexible, natural, and expressive speech synthesis.
+Unlike traditional methods, neural TTS leverages DNNs to model complex relationships between input text and speech, facilitating nuanced control over various speech characteristics.
+Early neural TTS systems, such as WaveNet [73] and Tacotron [74], laid the groundwork for controllability.
 1) Controlling prosody features like rhythm and intonation is vital for generating expressive and contextually appropriate speech.
-Neural-based TTS models achieve prosody control through explicit conditioning or learned latent representations [75, 15, 76, 77, 78].
-2) Speaker control has also gained significant improvement in neural-based TTS through speaker embeddings or adaptation techniques [79, 80, 81, 82].
-3) Besides, emotionally controllable TTS [83, 31, 20, 22, 32] has become a hot topic due to the strong modeling capability of DNNs, enabling the synthesis of speech with specific emotional tones such as happiness, sadness, anger, or neutrality.
+Neural TTS models achieve prosody control through explicit conditioning or learned latent representations [75, 15, 76, 77, 78].
+2) Speaker control has also gained significant improvement in neural TTS through speaker embeddings or adaptation techniques [79, 80, 81, 82].
+3) Besides, emotionally controllable TTS [83, 31, 20, 22, 32] has become a hot topic due to the strong modeling capability of DNNs, enabling the synthesis of speech with specific emotional tones, such as happiness, sadness, anger, or neutrality.
 These systems go beyond producing intelligible and natural-sounding speech, focusing on generating expressive output that aligns with the intended emotional context.
-4) Neural-based TTS can also manipulate timbre (vocal quality) [78, 84, 85, 14, 86, 87] and style (speech mannerisms) [88, 89, 90], allowing for creative and personalized applications.
+4) Neural TTS can also manipulate timbre (vocal quality) [78, 84, 85, 14, 86, 87] and style (speech mannerisms) [88, 89, 90], allowing for creative and personalized applications.
 These techniques lead to one of the most popular research topics, i.e., zero-shot TTS (particularly voice cloning) [82, 78, 91, 92].
@@
 These methods can emphasize or de-emphasize specific words or adjust the pronunciation of phonemes through speech editing or generation techniques.
-Neural-based TTS technologies represent a significant leap in the flexibility and quality of speech synthesis.
+Neural TTS technologies represent a significant leap in the flexibility and quality of speech synthesis.
 From prosody and emotion to speaker identity and style, these systems empower diverse applications in fields such as entertainment, accessibility, and human-computer interaction.
 LLM-based Synthesis.
-Here we pay special attention to LLM-based synthesis methods due to their superior context modeling capabilities compared to other neural-based TTS methods.
-LLMs, such as GPT [97, 98], T5 [99], and PaLM [100], have revolutionized various natural language processing (NLP) tasks with their ability to generate coherent, context-aware text.
+Here, we pay special attention to LLM-based synthesis methods due to their superior context modeling capabilities compared to other neural TTS methods.
+LLMs, such as generative pre-trained transformer (GPT) [97, 98], T5 [99], and pathways language model (PaLM) [100], have revolutionized various natural language processing (NLP) tasks with their ability to generate coherent, context-aware text.
 Recently, their utility has expanded into controllable TTS technologies [101, 102, 103, 104, 17].
@@
 Specifically, an LLM can detect emotional intent in sentences (e.g., “I’m thrilled” → happiness, “This is unfortunate” → sadness).
-The detected emotion is encoded as an auxiliary input to the TTS model, enabling modulation of acoustic features like prosody, pitch, and energy to align with the expressed sentiment.
-By leveraging LLMs’ capabilities in understanding and generating rich contextual information, these systems can achieve enhanced and fine-grained control over various speech attributes such as prosody, emotion, style, and speaker characteristics [105, 31, 106].
-Integrating LLMs into TTS systems represents a significant step forward, enabling more dynamic and expressive speech synthesis.
+The detected emotion is encoded as an auxiliary input to the TTS model, enabling the modulation of acoustic features like prosody, pitch, and energy to align with the expressed sentiment.
+By leveraging LLMs’ capabilities in understanding and generating rich contextual information, these systems can achieve enhanced and fine-grained control over various speech attributes, such as prosody, emotion, style, and speaker characteristics [105, 31, 106].
+Integrating LLMs into TTS represents a significant step forward, enabling more dynamic and expressive speech synthesis.
 I-C Organization of This Survey
-This paper first presents a comprehensive and systematic review of controllable TTS technologies, with a particular focus on model architectures, control methodologies, and feature representations.
+This survey first presents a comprehensive and systematic review of controllable TTS technologies, with a particular focus on model architectures, control strategies, and feature representations.
 To establish a foundational understanding, this survey begins with an introduction to the TTS pipeline in Section II.
@@
 Section V presents a comprehensive review of datasets and evaluation metrics.
-Section VI provides an in-depth analysis of the challenges encountered in achieving controllable TTS systems and discusses future directions.
-Section VII explores the broader impacts of controllable TTS technologies and identifies promising future research directions, followed by the conclusion in Section VIII.
+Section VI provides an in-depth analysis of the challenges encountered in achieving controllable TTS systems and identifies promising future research directions.
+Section VII explores the broader impacts of controllable TTS technologies, followed by the conclusion in Section VIII.
 II TTS Pipeline
-Figure 2: The general pipeline of controllable TTS from the perspective of network structure.
-Linguistic analysis is necessary for parametric and a few neural-based methods but is no longer needed for most modern neural-based methods.
-In this paper, we only review neural-based controllable TTS methods and do not investigate acoustic features (e.g., MCC [107], LSP [108], F0 [109]) used in early TTS methods.
+Figure 2: General pipeline of controllable TTS from the perspective of network structure.
+Linguistic analysis is necessary for parametric and a few neural methods but is no longer needed for most modern neural methods.
+In this paper, we only review neural model-based controllable TTS methods and do not investigate acoustic features (e.g., MFCC [107], LSP [108], F0 [109]) used in early TTS methods.
 In this section, we elaborate on the general pipeline that supports controllable TTS technologies, including acoustic models, speech vocoders, and feature representations.
-Fig.2 depicts the general pipeline of controllable TTS, containing various model architectures and feature representations, but the control strategies will be discussed in Section IV.
-Readers can jump to Section III if familiar with TTS pipelines.
+Fig.
+2 depicts the general pipeline of controllable TTS, containing various model architectures and feature representations, but the control strategies will be discussed in Section IV.
+Readers familiar with TTS pipelines may skip ahead to Section III.
 II-A Overview
-A TTS pipeline generally contains three key components, i.e., linguistic analyzer, acoustic model, speech vocoder, and with a conditional input, e.g., prompts, for controllable speech synthesis.
-Besides, some end-to-end methods use a single model to encode the input and decode the speech waveforms without generating intermediate features like mel-spectrograms [110].
-Linguistic analyzer aims to extract linguistic features, e.g., phoneme duration and position, syllable stress, and utterance level, from the input text, which is a necessary step in HHM-based methods [64, 65] and a few neural-based methods [111, 112], but is time-consuming and error-prone.
+A TTS pipeline generally contains three key components, i.e., linguistic analyzer, acoustic model, and speech vocoder, where a conditional input, e.g., prompts, can be processed for controllable speech synthesis.
+Linguistic analyzer aims to extract linguistic features, e.g., phoneme duration and position, syllable stress, and utterance level, from the input text, which is a necessary step in HMM-based methods [64, 65] and a few neural model-based methods [110, 111], but is time-consuming and error-prone.
 Acoustic model is a parametric or neural model that predicts the acoustic features from the input texts.
-Modern neural-based acoustic models like Tacotron [74] and later works [15, 76, 113] directly take character [114] or word embeddings [115] as the input, which is much more efficient than previous methods.
+Modern neural acoustic models like Tacotron [74] and later works [15, 76, 112] directly take character [113] or word embeddings [114] as the input, which is much more efficient than previous methods.
 Speech vocoder is the last component that converts the intermediate acoustic features into a waveform that can be played back.
 This step bridges the gap between the acoustic features and the actual sounds produced, helping to generate high-quality, natural-sounding speech [73, 115].
+Besides, some end-to-end methods use a single model to encode the input and decode the speech waveforms without generating intermediate features like mel-spectrograms.
 Tan et al.
@@
 II-B Acoustic Models
-Acoustic modeling is a crucial step in TTS because it ensures the generated acoustic features capture the subtleties of human speech.
+Acoustic modeling is a crucial step in TTS because it ensures that the generated acoustic features capture the subtleties of human speech.
 By accurately modeling acoustic features, modern TTS systems can help generate high-quality and expressive audio that sounds close to human speech.
@@
 RNN-based Models.
-Recurrent Neural Networks (RNNs) proved particularly effective in early neural-based TTS due to their ability to model sequential data and long-range dependencies, which helps in capturing the sequential nature of speech, such as the duration and natural flow of phonemes.
+Recurrent Neural Networks (RNNs) proved particularly effective in early neural TTS due to their ability to model sequential data and long-range dependencies, which helps in capturing the sequential nature of speech, such as the duration and natural flow of phonemes.
 Typically, these models have an encoder-decoder architecture, where an encoder encodes input linguistic features, such as phonemes or text, into a fixed-dimensional representation, and the decoder sequentially decodes this representation into acoustic features (e.g., mel-spectrogram frames) that capture the frequency and amplitude of sound over time.
-Tacotron 2 [75] is one of the pioneering TTS models that uses RNNs with an attention mechanism, which helps align the text sequence with the generated acoustic features.
+Tacotron 2 [75] is one of the pioneering TTS models that use RNNs with an attention mechanism, which helps align the text sequence with the generated acoustic features.
 It takes raw characters as input and produces mel-spectrogram frames, which are subsequently converted to waveforms.
@@
 Unlike RNNs, which process sequential data frame by frame, CNNs process the entire sequence at once by applying filters across the input texts.
-This parallel approach enables faster training and inference, making CNN-based TTS particularly appealing for real-time and low-latency applications.
+This parallel approach enables faster training and inference, making CNN-based TTS particularly appealing for real-time and low-latency applications [73, 16, 120, 121].
 Furthermore, by stacking multiple convolutional layers with varying kernel sizes or dilation rates, CNNs can capture both short-range and long-range dependencies, which are essential for natural-sounding speech synthesis.
-Deep Voice [122] is one of the first prominent CNN-based TTS models by Baidu, designed to generate mel-spectrograms directly from phoneme or character input.
-ParaNet [123] also utilizes a RNN model to achieve sequence-to-sequence mel-spectrogram generation.
+Deep Voice [16] is one of the first prominent CNN-based TTS methods, designed to generate mel-spectrograms directly from phoneme or character input.
+ParaNet [122] also utilizes a CNN model to achieve sequence-to-sequence mel-spectrogram generation.
 It uses a non-autoregressive architecture, which enables significantly faster inference by predicting multiple time steps simultaneously.
 Transformer-based Models.
-Transformer model [124] uses self-attention layers to capture relationships within the input sequence, making them well-suited for tasks requiring an understanding of global contexts, such as prosody and rhythm in TTS.
+The transformer model [123] uses self-attention layers to capture relationships within the input sequence, making them well-suited for tasks requiring an understanding of global contexts, such as prosody and rhythm in TTS.
 Transformer-based TTS models often employ an encoder-decoder architecture, where the encoder processes linguistic information (e.g., phonemes or text) and captures contextual relationships, and the decoder generates acoustic features (like mel-spectrograms) from these encoded representations, later converted to waveforms by a vocoder.
@@
 Other Acoustic Models.
-In TTS, GANs [127, 128, 129], VAEs [130, 18], and diffusion models [113, 131] can also be used as acoustic models.
-Flow-based methods [132, 133] are also popular in waveform generation.
+In TTS, generative adversarial networks (GANs) [126, 127, 128], variational autoencoders (VAEs) [129, 18], and diffusion models [112, 130] can also be used as acoustic models.
+Flow-based methods [131, 132] are also popular in acoustic feature generation.
+Recently proposed flow-based controllable TTS methods will be discussed in Section IV.
 Refer to the survey paper from Tan et al.
@@
 II-C Speech Vocoders
-Vocoders are essential for converting acoustic features such as mel-spectrograms into intelligible audio waveforms and are vital in determining the naturalness and quality of synthesized speech.
+Vocoders are essential for converting acoustic features, such as mel-spectrograms, into intelligible audio waveforms and are vital in determining the naturalness and quality of synthesized speech.
 We broadly categorize existing vocoders according to their model architectures, i.e., RNN-, CNN-, GAN-, and diffusion-based vocoders.
@@
 CNN-based Vocoders.
-By leveraging the parallel nature of convolutional operations, CNN-based vocoders [73, 140, 141] can generate high-quality speech more efficiently, making them ideal for real-time applications.
-A key strength of CNN-based vocoders is their ability to balance synthesis quality and efficiency.
+By leveraging the parallelism of convolutional operations, CNN-based vocoders [73, 139, 140] can achieve higher speech quality and more efficient synthesis compared to parametric vocoders [133, 134], making them ideal for applications that demand real-time and natural speech synthesis.
 However, they often require extensive training data and careful hyperparameter tuning to achieve optimal performance.
-WaveNet [73] is a probabilistic autoregressive model that generates waveforms sample by sample conditioned on all preceding samples and auxiliary inputs, such as linguistic features and mel-spectrograms.
+WaveNet [73] is a probabilistic autoregressive model that generates waveforms sample by sample, conditioned on all preceding samples and auxiliary inputs, such as linguistic features and mel-spectrograms.
 It employs stacks of dilated causal convolutions, enabling long-range dependence modeling in speech signals without relying on recurrent connections.
@@
 A key advantage of GAN-based vocoders is their parallel inference capability, enabling real-time synthesis with lower computational costs compared to autoregressive models.
-However, training GANs can be challenging due to instability and mode collapse.
-Despite these challenges, GAN-based vocoders continue to advance the state-of-the-art in neural vocoding, offering a compelling combination of speed and audio quality.
+However, training GANs can be challenging due to instability and mode collapse caused by imbalanced adversarial dynamics, vanishing gradients, and the generator overfitting to limited patterns [145, 146].
+Despite these challenges, GAN-based vocoders continue to advance the SOTA in neural vocoding, offering a compelling combination of speed and audio quality.
 Diffusion-based Vocoders.
-Inspired by diffusion probabilistic models [146] that have shown success in visual generation tasks, diffusion-based vocoders [147, 113, 148, 149, 150] present a novel approach to natural-sounding speech synthesis.
+Inspired by diffusion probabilistic models [147] that have shown success in visual generation tasks, diffusion-based vocoders [148, 112, 149, 150, 151] present an alternative approach to natural-sounding speech synthesis.
 The core mechanism of diffusion-based vocoders involves two stages: a forward process and a reverse process.
@@
 Other Vocoders.
-There are also many other types of vocoders such as flow-based [152, 153, 154, 155, 156] and VAE-based vocoders [157, 158, 159].
-These methods provide unique strengths for speech synthesis such as efficiency and greater flexibility in modeling complex speech variations.
+There are also many other types of vocoders, such as flow-based [153, 154, 155, 156, 157] and VAE-based vocoders [122, 158, 159].
+These methods provide unique strengths for speech synthesis, such as efficiency and greater flexibility in modeling complex speech variations.
 Readers can refer to the survey paper from Tan et al.
@@
 Fully end-to-end TTS methods [160, 161, 76, 159, 162] directly generate speech waveforms from textual input, simplifying the “acoustic model → vocoder” pipeline and achieving efficient speech generation.
-Char2Wav [160] is an early neural text-to-speech (TTS) system that directly synthesizes speech waveforms from character-level text input.
+Char2Wav [160] is an early neural TTS system that directly synthesizes speech waveforms from character-level text input.
 It integrates two components and jointly trains them: a recurrent sequence-to-sequence model with attention, which predicts acoustic features (e.g., mel-spectrograms) from text, and a SampleRNN-based neural vocoder [135] that generates waveforms from these features.
@@
 VITS [159] is another fully end-to-end TTS framework.
-It integrates a variational autoencoder (VAE) with normalizing flows [163] and adversarial training, enabling the model to learn latent representations that capture the intricate variations in speech, such as prosody and style.
+It integrates a VAE with normalizing flows [163] and adversarial training, enabling the model to learn latent representations that capture the intricate variations in speech, such as prosody and style.
 VITS combines non-autoregressive synthesis with stochastic latent variable modeling, achieving real-time waveform generation without compromising naturalness.
-There are more end-to-end TTS models such as Tacotron [74], ClariNet [161], and EATS [162], refer to another survey [42] for more details.
+There are more end-to-end TTS models such as Tacotron [74], ClariNet [161], and EATS [162], readers can refer to another survey [42] for more details.
 End-to-end controllable methods that emerged in recent years will be discussed in Section IV.
-TABLE II: Popular open-source speech quantization methods.
-SSCP: Self-supervised context prediction, RVQ: Residual vector quantization [168].
+TABLE II: Popular open-source speech quantization and tokenization methods.
+SSCP: Self-supervised context (token) prediction, RVQ: Residual vector quantization [170].
 II-E Acoustic Feature Representations
 In TTS, the choice of acoustic feature representations impacts the model’s flexibility, quality, expressiveness, and controllability.
-This subsection investigates continuous representations and discrete tokens as shown in Fig.2, along with their pros and cons for TTS applications.
+This subsection investigates continuous representations and discrete tokens as shown in Fig.
+2, along with their pros and cons for TTS applications.
 Continuous Representations.
-Continuous representations (e.g., mel-spectrograms) of intermediate acoustic features use a continuous feature space to represent speech signals.
+Continuous representations (e.g., mel-spectrograms and VAE features) of intermediate acoustic features use a continuous feature space to model speech signals.
 These representations often involve acoustic features that capture frequency, pitch, and other characteristics without discretizing the signal.
-The advantages of continuous features are: 1) Continuous representations retain fine-grained detail, enabling more expressive and natural-sounding speech synthesis.
+The advantages of continuous features are: 1) Continuous representations retain fine-grained detail, enabling expressive and natural-sounding speech synthesis.
 2) Since continuous features inherently capture variations in tone, pitch, and emphasis, they are well-suited for prosody control and emotional TTS.
-3) Continuous representations are more robust to information loss and can avoid quantization artifacts, allowing smoother, less distorted audio.
+3) Continuous representations are robust to information loss and can avoid quantization artifacts, allowing reconstruction of smooth audio.
 GAN-based [115, 143, 144] and diffusion-based methods [149, 148] often utilize continuous feature representations, i.e., mel-spectrograms.
-However, continuous representations are typically more computationally demanding and require larger models and memory, especially in high-resolution audio synthesis.
+However, continuous representations are typically computationally demanding and require large-scale models and training datasets, especially in high-resolution audio synthesis.
 Discrete Tokens.
 In discrete token-based TTS, the intermediate acoustic features (e.g., quantized units or phoneme-like tokens) are discrete values, similar to words or phonemes in languages.
-These are often produced using quantization techniques or learned embeddings, such as HuBERT [166] and SoundStream [168].
-The advantages of discrete tokens are: 1) Discrete tokens can encode phonemes or sub-word units, making them concise and less computationally demanding to handle.
-2) Discrete tokens often allow TTS systems to require fewer samples to learn and generalize, as the representations are compact and simplified.
-3) Using discrete tokens simplifies cross-modal TTS applications like voice cloning or translation-based TTS, as they map well to text-like representations such as LLM tokens.
+These are often produced using quantization techniques or learned token embeddings, such as HuBERT [166] and SoundStream [170].
+The advantages of discrete tokens are: 1) Discrete tokens can encode phonemes or sub-word units, making them concise and computationally efficient to handle.
+2) Discrete tokens often allow TTS systems to require fewer samples to learn and generalize, as compared with continuous representations, since the representations are compact and simplified.
+3) Using discrete tokens simplifies cross-modal TTS applications like voice cloning or text prompt-based TTS, as they map well to text-like representations such as LLM tokens.
 LLM-based [78, 103, 106, 105] and zero-shot TTS methods [17, 78, 87] often adopt discrete tokens as their acoustic features.
 However, discrete representation learning may result in information loss or lack the nuanced details that can be captured in continuous representations.
-Table IV and III summarize the types of acoustic features of representative methods. Table II summarizes popular open-source speech quantization methods.
+Speech Quantization vs.
+Tokenization.
+It is worth noting that quantization and tokenization serve distinct purposes in speech processing.
+Quantization is primarily used for high-fidelity compression, reducing the precision of numerical representations (e.g., from 32-bit floating point to 8-bit integers) while preserving model performance.
+In speech synthesis, quantization is often used in waveform generation (e.g., codec-based approaches like EnCodec [171]) and neural vocoders to compress audio signals without significant loss of perceptual quality.
+Tokenization, on the other hand, is a discretization process that segments continuous data into meaningful units.
+In speech tasks, tokenization extracts semantically relevant representations such as phonemes, characters, or learned speech units (e.g., HuBERT [166] and Wav2Vec 2.0 [165]).
+This makes tokenization particularly suitable for speech-to-text (ASR), TTS, and multimodal NLP tasks, where aligning speech with textual information is crucial.
+Tokenization also facilitates training language models on speech data by enabling linguistic or learned unit-based processing rather than raw audio waveform modeling.
+Table II summarizes popular open-source speech quantization and tokenization methods.
+Tables IV and III summarize the types of acoustic features of representative methods.
 III Uncontrollable TTS
@@
 This shift highlights the integration of both local and global information to produce speech with human-like quality and naturalness.
-This survey explores UC-TTS evolution, emphasizing the role of local and global information in enhancing speech fidelity and expressiveness.
-In the context of UC-TTS, ”uncontrollable” refers to the absence of explicit control mechanisms for speech features such as emotion, timbre, and speaker style.
-Despite this lack of explicit control, the goal is to achieve natural, fluid speech while minimizing issues like mispronunciations and omissions.
+This section briefly investigates UC-TTS methods, emphasizing the role of local and global information in enhancing speech fidelity and expressiveness.
+In the context of UC-TTS, the term “uncontrollable” refers to the absence of explicit control mechanisms for speech features such as emotion, timbre, and speaking style.
+Despite this, the goal is to achieve natural, fluid speech while mitigating issues like mispronunciations and omissions.
 III-A Early Approaches: Statistical Models
-Early Text-To-Speech (TTS) systems relied on statistical models such as Hidden Markov Models (HMMs) [64, 65] and early neural network-based parametric methods [111, 112].
+Early TTS systems relied on statistical models such as HMMs [64, 65] and early neural network-based parametric methods [110, 111].
 These models operated at the frame level, using acoustic models and vocoders for text-to-speech conversion.
@@
 III-B Sequence-to-Sequence Models
-The emergence of sequence-to-sequence models represents a significant breakthrough by removing the need for explicit linguistic features, thereby enabling the capture of the nuances and idiosyncrasies of human speech.
-Models such as Tacotron [74] and Tacotron 2  [175] utilize recurrent neural networks (RNNs) with attention mechanisms to effectively model the complex, nonlinear nature of speech sequences.
-These innovations allow for precise tuning of speech parameters, enhancing prosody and rhythm by modeling entire utterances rather than isolated phonetic units.
+The emergence of sequence-to-sequence models represents a significant breakthrough by removing the need for explicit linguistic features, thereby enabling the capture of the nuances of human speech.
+Models such as Tacotron [74] and Tacotron 2 [75] utilize RNNs with attention mechanisms to effectively model the complex, nonlinear nature of speech sequences.
+These innovations enable the tuning of speech parameters, enhancing prosody and rhythm by modeling entire utterances instead of isolated phonetic units.
 Building on these advancements, Deep Voice 3[121] introduces a fully convolutional sequence-to-sequence architecture that significantly accelerates training speed compared to RNN-based models.
 This approach achieves training times an order of magnitude faster, enabling scalability to handle large datasets.
-Additionally, the use of a position-augmented attention mechanism in Deep Voice 3 enhances the naturalness of synthesized speech, achieving competitive mean opinion scores, especially when paired with advanced neural vocoders like WaveNet.
-This development not only improves training efficiency but also enhances the scalability and naturalness of text-to-speech systems.
+Additionally, the position-augmented attention mechanism in Deep Voice 3 enhances the naturalness of synthesized speech, achieving competitive mean opinion scores, especially when paired with advanced neural vocoders like WaveNet.
+This development not only improves training efficiency but also enhances the scalability and naturalness of TTS systems.
 III-C Transformer-based Models
@@
 To address these, techniques such as relative position encodings and localized attention were integrated [123].
-III-D Advanced Architectures: Integrating Flow and Diffusion Models
-Recent advancements have shifted towards integrating global information within end-to-end architectures to enhance speech naturalness and coherence.
+III-D Integrating Flow and Diffusion Models
+Recent advancements have shifted toward integrating global information within end-to-end architectures to enhance speech naturalness and coherence.
 Flow-based models like Glow-TTS[132] and Flow-TTS[131] exemplify this by employing invertible transformations that maintain the balance between local precision and global coherence.
 These architectures enable the synthesis of high-fidelity speech by modeling complex dependencies across the entire utterance, thus improving the overall fluidity and naturalness of the generated speech.
-Moreover, the introduction of diffusion models in TTS, such as WaveGrad 2[178], highlights the shift towards models that can iteratively refine speech output.
-These models use score matching and diffusion processes to generate speech directly from phoneme sequences, effectively capturing both local nuances and overarching global patterns.
+Moreover, the introduction of diffusion models in TTS, such as WaveGrad 2[178], highlights the shift toward models that can iteratively refine speech output.
+These models use score-matching and diffusion processes to generate speech directly from phoneme sequences, effectively capturing both local nuances and overarching global patterns.
 The iterative nature of these models allows for adjustments that enhance the quality of the synthesized audio, accommodating variations in speech without explicit control over specific attributes.
-The integration of adversarial training and variational autoencoders (VAEs) further exemplifies the evolution towards incorporating global information.
+The integration of adversarial training and VAEs further exemplifies the evolution toward incorporating global information.
 Systems like VITS[159] leverage these techniques to enhance expressiveness and naturalness by learning complex mappings between text and speech.
-This approach allows the model to manage variations in prosody and rhythm inherently derived from the textual input, aligning with the objectives of UC-TTS to produce diverse and natural speech outputs.
+This approach allows the model to manage variations in prosody and rhythm that are inherently derived from the textual input, aligning with the objectives of UC-TTS to produce diverse and natural speech.
 The evolution from HMMs to advanced architectures in UC-TTS exemplifies progress toward synthesizing speech that is both expressive and precise.
@@
 IV Controllable TTS
-In this section, we first review recent TTS work from the perspective of model architecture, followed by a detailed discussion of control modes in controllable TTS.
-IV-A Model Architectures
+In this section, we first review recent TTS work from the perspective of model architecture, followed by a detailed discussion of control strategies in controllable TTS, which is the core part of this survey.
 Current model architectures can be broadly classified into two main categories: The first is the non-autoregressive (NAR) generative models, which are based on HMMs, neural networks, VAEs, diffusion models, flow matching, and other NAR techniques.
@@
 We summarize the NAR-based and AR-based controllable TTS methods in Table III and Table IV, respectively.
-IV-A1 Fully Non-Autoregressive (NAR) Architectures
-HMM-based Approaches.
-In the realm of Controllable Text-To-Speech (CTTS), advancements in Hidden Markov Model (HMM) architectures have significantly enhanced the manipulation of speech elements such as emotion and prosody.
-Yamagishi et al.
-[70] pioneered this field by introducing style-dependent and style-mixed modeling, which allowed precise emulation of human-like emotional nuances and versatile synthesis across various styles by incorporating style as a contextual variable.
-Building on this foundation, Qin et al.[179] developed the ”average emotion model,” which utilized MLLR-based adaptation to modulate emotions like happiness and sadness even with limited data, thus advancing the emotional intelligence of synthetic speech systems.
-Furthering expressive variability, Nose et al.[119] integrated subjective style intensities and a multiple-regression global variance model into HMM frameworks, addressing over-smoothing and enabling nuanced emotional expressions.
-Lorenzo-Trueba et al.[72] expanded on these capabilities with CSMAPLR adaptation, introducing ”emotion transplantation” to transfer emotional states between speakers while preserving voice distinctiveness, enhancing personalized human-computer interaction.
-These innovations in HMM architectures have broadened the expressiveness and individuality in synthetic speech, augmenting technological interfaces and paving the way for future developments in adaptive, lifelike speech solutions.
+TABLE III: A summary of existing non-autoregressive controllable neural-based methods.
+Method Zero-shot TTS Controlability Model Architectures Acoustic Feature Release Time
+Pit. Ene. Spe. Pro. Tim. Emo. Env. Des. Acoustic Model Vocoder
+FastSpeech[15] Transformer WaveGlow [179] MelS 2019.05
+FastSpeech 2[180] Transformer Parallel WaveGAN [143] MelS 2020.06
+FastPitch[77] Transformer WaveGlow MelS 2020.06
+Parallel Tacotron[181] Transformer + CNN WaveRNN [136] MelS 2020.10
+StyleTagging-TTS [182] Transformer + CNN HiFi-GAN [115] MelS 2021.04
+SC-GlowTTS [183] Transformer + Flow HiFi-GAN MelS 2021.06
+Meta-StyleSpeech [184] Transformer MelGAN [144] MelS 2021.06
+DelightfulTTS [185] Transformer + CNN HiFiNet [185] MelS 2021.11
+YourTTS [82] Transformer + Flow HiFi-GAN LinS 2021.12
+StyleTTS [88] CNN + RNN HiFi-GAN MelS 2022.05
+GenerSpeech [90] Transformer + Flow HiFi-GAN MelS 2022.05
+Cauliflow[186] BERT + Flow UP WaveNet [187] MelS 2022.06
+CLONE[188] Transformer + CNN WaveNet [73] MelS + LinS 2022.07
+PromptTTS [101] BERT + Transformer HiFi-GAN MelS 2022.11
+Grad-StyleSpeech [189] Score-based Diffusion HiFi-GAN MelS 2022.11
+NaturalSpeech 2 [86] Diffusion RVQ-based Codec [86] Latent Feature 2023.04
+PromptStyle [190] VITS + Flow HiFi-GAN MelS 2023.05
+StyleTTS 2 [89] Flow-based Diffusion + GAN HifiGAN / iSTFTNet [191] MelS 2023.06
+VoiceBox [192] Transformer + Flow HiFi-GAN MelS 2023.06
+MegaTTS 2 [193] Decoder-only Transformer + GAN HiFi-GAN MelS 2023.07
+PromptTTS 2 [102] Diffusion RVQ-based Codec Latent Feature 2023.09
+VoiceLDM [194] Diffusion HiFi-GAN MelS 2023.09
+DurIAN-E[195] CNN + RNN HiFi-GAN MelS 2023.09
+PromptTTS++ [104] Transformer + Diffusion BigVGAN [196] MelS 2023.09
+SpeechFlow [197] Transformer + Flow HiFi-GAN MelS 2023.10
+P-Flow [198] Transformer + Flow HiFi-GAN MelS 2023.10
+E3 TTS [199] Diffusion Not required Waveform 2023.11
+HierSpeech++ [200] Transformer + VAE + Flow BigVGAN MelS 2023.11
+Audiobox [201] Transformer + Flow EnCodec [171] MelS 2023.12
+FlashSpeech [202] Latent Consistency Model EnCodec Token 2024.04
+NaturalSpeech 3 [87] Transformer + Diffusion FACodec [87] Token 2024.04
+InstructTTS [105] Transformer + Diffusion HiFi-GAN Token 2024.05
+ControlSpeech [106] Transformer + Diffusion FACodec Token 2024.06
+AST-LDM [203] Diffusion + VAE HiFi-GAN MelS 2024.06
+SimpleSpeech [204] Transformer + Diffusion SQ Codec [204] Token 2024.06
+DiTTo-TTS [205] DiT + VAE BigVGAN MelS 2024.06
+E2 TTS [206] Transformer + Flow BigVGAN MelS 2024.06
+MobileSpeech [207] Transformer Vocos [208] Token 2024.06
+DEX-TTS [209] Diffusion HiFi-GAN MelS 2024.06
+ArtSpeech [210] RNN + CNN HiFI-GAN MelS + Energy + F0 2024.07
+CCSP [211] Diffusion RVQ-based Codec [211] Token 2024.07
+SimpleSpeech 2 [212] Flow-based DiT SQ Codec Token 2024.08
+E1 TTS [213] DiT + Flow BigVGAN Token + MelS 2024.09
+StyleTTS-ZS [214] Flow-based Diffusion + GAN Mel-based Decoder [214] MelS 2024.09
+NansyTTS [215] Transformer NANSY++ [215] MelS 2024.09
+NanoVoice [216] Diffusion BigVGAN MelS 2024.09
+MS2KU-VTTS [217] Transformer BigVGAN MelS 2024.10
+MaskGCT [78] Transformer + Flow Vocos Token 2024.10
+EmoSphere++ [218] Transformer + Flow BigVGAN MelS 2024.11
+EmoDubber [219] Transformer + Flow Flow-based Vocoder [219] MelS 2024.12
+HED [220] Flow-based Diffusion Vocos MelS 2024.12
+DiffStyleTTS [221] Transformer + Diffusion HiFi-GAN MelS 2025.01
+DrawSpeech [222] Diffusion HiFi-GAN MelS 2025.01
+ProEmo [223] Transformer HiFi-GAN MelS 2025.01
+Abbreviations: Pit(ch), Ene(rgy), Spe(ed), Pro(sody), Tim(bre), Emo(tion), Env(ironment), Des(cription). MelS: Mel Spectrogram. LinS: Linear Spectrogram.
+IV-A Non-Autoregressive Architectures
+In non-autoregressive TTS models, the model generates the entire output sequence $\mathbf{y}=(y_{1},y_{2},\dots,y_{T})$ at once, conditioned on the input sequence $\mathbf{x}=(x_{1},x_{2},\dots,x_{T})$.
+The probability distribution for generating the sequence is:
+$P(\mathbf{y}|\mathbf{x})=P(\mathbf{y}|\mathbf{x},\theta),$ (1)
+Where $P(\mathbf{y}|\mathbf{x})$ is the likelihood of the output sequence $\mathbf{y}$ given the input $\mathbf{x}$, and $\theta$ represents the parameters of the model (e.g., weights).
+Since the output sequence is predicted simultaneously, the model learns to capture dependencies in a way that does not rely on previously generated outputs.
 Transformer-based Approaches.
-Advancements in Controllable Text-to-Speech (TTS) technology highlight the integration of deep learning with audio processing, driven by Transformer-based architectures.
-Ren et al.[15] introduced FastSpeech, a feed-forward non-autoregressive Transformer model that significantly enhances TTS efficiency by reducing inference time and improving the stability issues found in autoregressive models like Tacotron 2.
+Advancements in controllable TTS technology highlight the integration of deep learning with audio processing, driven by Transformer-based architectures.
+Ren et al.
+[15] introduced FastSpeech, a feed-forward non-autoregressive Transformer model that significantly enhances TTS efficiency by reducing inference time and improving the stability issues found in autoregressive models like Tacotron 2.
 This model provides precise control over prosodic features through duration prediction, effectively tackling the one-to-many mapping challenge.
@@
 VAE-based Approaches.
-Recent advancements in Controllable Text-To-Speech (TTS) systems are largely driven by the integration of Variational Autoencoder (VAE) architectures, which enhance the flexibility and precision of speech modulation.
-Zhang et al.[18] pioneered the use of VAEs in end-to-end speech synthesis, creating disentangled latent representations that allow effective style control and transfer, especially in prosody and emotion management, outperforming the Global Style Token model in style transfer tasks.
+Recent advancements in controllable TTS systems are largely driven by the integration of VAE architectures, which enhance the flexibility and precision of speech modulation.
+Zhang et al.[18] pioneered the use of VAEs in end-to-end speech synthesis, creating disentangled latent representations that allow effective style control and transfer, especially in prosody and emotion management, outperforming the Global Style Token (GST) model [19] in style transfer tasks.
 Building on this, Hsu et al.[129] developed a hierarchical generative model with a conditional VAE framework and a Gaussian mixture model, enabling precise control over complex speech attributes such as environment and style, thus improving expressive speech synthesis through refined noise and speaker characteristic management.
 Liu et al.[188] further advanced the field with the CLONE model, a single-stage TTS system that resolves the one-to-many mapping issue and enhances high-frequency information reconstruction.
-By employing a conditional VAE with normalizing flows and a dual path adversarial training mechanism with multi-band discriminators, CLONE achieves nuanced control over prosody and energy, demonstrating superior performance in both speech quality and prosody control compared to state-of-the-art models.
+By employing a conditional VAE with normalizing flows and a dual-path adversarial training mechanism with multi-band discriminators, CLONE achieves nuanced control over prosody and energy, demonstrating superior performance in both speech quality and prosody control compared to state-of-the-art models.
 These collective innovations highlight the adaptability of VAEs in managing complex speech generation tasks, marking significant progress toward more dynamic and versatile TTS technologies, with ongoing research promising even greater advancements.
@@
 NaturalSpeech 3 [87] decomposes speech into distinct subspaces that represent different attributes and generates each subspace independently.
-DEX-TTS [183] improves DiT-based diffusion networks by applying overlapping patchify and convolution-frequency patch embedding strategies.
-E3 TTS [184] models the temporal structure of the waveform through the diffusion process, eliminating the need for any intermediate representations such as spectrogram features or alignment information.
+DEX-TTS [209] improves DiT-based diffusion networks by applying overlapping patching and convolution-frequency patch embedding strategies.
+E3 TTS [199] models the temporal structure of the waveform through the diffusion process, eliminating the need for any intermediate representations, such as spectrogram features or alignment information.
 Applying diffusion models to TTS requires a complex pipeline due to the need for precise temporal alignment between text and speech and the high fidelity required for audio data.
@@
 Similarly, SimpleSpeech [204] proposes a speech codec model (SQ-Codec) based on scalar quantization and uses the sentence duration to control the generated speech length.
+Besides TTS, some text-to-audio models can also perform TTS and are worth referring to, such as AudioLDM [224], AudioLDM2 [225], Make-An-Audio [226], and CosyAudio [227].
 Flow-based Approaches.
-Flow-based methods leverage invertible flow transformations to learn mappings from target speech features to simple distributions, typically standard Gaussian distributions.
+Flow-based methods leverage invertible flow transformations [163, 228] to learn mappings from target speech features to simple distributions [179], typically standard Gaussian distributions.
 Due to their invertibility, this mechanism can directly sample from the simple distribution and generate high-fidelity speech in the reverse direction.
-Audiobox [187] and P-flow [188] employ non-autoregressive flow-matching models for efficient and stable speech synthesis.
-VoiceBox [189] also employs a flow-matching to generate speech, effectively casting the TTS task into a speech infilling task.
+Audiobox [201] and P-flow [198] employ non-autoregressive flow-matching models [228] for efficient and stable speech synthesis.
+VoiceBox [192] also employs flow-matching to generate speech, effectively casting the TTS task into a speech infilling task.
 SpeechFlow [197] is trained on 60k hours of untranscribed speech with flow matching and mask conditions and can be fine-tuned with task-specific data to match or surpass existing expert models.
@@
 FlashSpeech [202] is built on a latent consistency model and applies a novel adversarial consistency training approach that can train from scratch without the need for a pre-trained diffusion model as the teacher, achieving speech generation in one or two steps.
-Recently, E2 TTS [193] converts text input into a character sequence with filler tokens and trains a mel spectrogram generator based on audio infilling task, achieving human-level naturalness.
+Recently, E2 TTS [206] converts text input into a character sequence with filler tokens and trains a mel spectrogram generator based on an audio infilling task, achieving human-level naturalness.
 Inspired by E2 TTS, F5-TTS [229] refines the text representation with ConvNext v2 [230], facilitating easier alignment with speech.
@@
 Other NAR Approaches.
-Other works leverage GAN-based or Masked Generative-based methods for TTS generation.
-StyleTTS 2 [89] employs large pre-trained speech language models (SLMs) such as Wav2Vec 2.0 [165], HuBERT [166], and WavLM [200] as discriminators, in combination with a novel differentiable duration modeling approach.
+Other works leverage GAN-based or masked generative model-based methods for TTS generation.
+StyleTTS 2 [89] employs large pre-trained speech language models (SLMs) such as Wav2Vec 2.0 [165], HuBERT [166], and WavLM [233] as discriminators in combination with a novel differentiable duration modeling approach.
 This setup uses SLM representations to enhance the naturalness of the synthesized speech.
-MaskGCT [78] proposes masked generative transformers without requiring text-speech alignment supervision and phone-level duration.
-The model employs a two-stage system, both trained using a mask-and-predict learning paradigm.
-IV-A2 Autoregressive (AR) Architectures
+MaskGCT [78] proposes masked generative transformers without requiring text-speech alignment supervision and phone-level duration prediction.
+The model employs a two-stage system trained using a mask-and-predict learning paradigm.
+TABLE IV: A summary of existing autoregressive controllable neural-based methods.
+Method Zero-shot TTS Controlability Model Architectures Acoustic Feature Release Time
+Pit. Ene. Spe. Pro. Tim. Emo. Env. Des. Acoustic Model Vocoder
+Prosody-Tacotron [234] RNN WaveNet MelS 2018.03
+GST-Tacotron [235] CNN + RNN Griffin-Lim LinS 2018.03
+GMVAE-Tacotron [129] VAE WaveRNN MelS 2018.12
+VAE-Tacotron [18] CNN + RNN WaveNet MelS 2019.02
+DurIAN [138] CNN + RNN Multi-band WaveRNN [138] MelS 2019.09
+Flowtron [236] CNN + RNN WaveGlow MelS 2020.07
+MsEmoTTS [83] CNN + RNN WaveRNN MelS 2022.01
+VALL-E [85] Decoder-only Transformer EnCodec Token 2023.01
+SpearTTS [237] Decoder-only Transformer SoundStream [170] Token 2023.02
+VALL-E X [238] Decoder-only Transformer EnCodec Token 2023.03
+Make-A-Voice [239] Encoder-decoder Transformer Unit-based Vocoder [239] Token 2023.05
+TorToise [240] Decoder-only Transformer + Diffusion UnivNet [241] MelS 2023.05
+MegaTTS [91] Decoder-only Transformer + GAN HiFi-GAN MelS 2023.06
+SC VALL-E [242] Decoder-only Transformer EnCodec Token 2023.07
+Salle [243] Decoder-only Transformer EnCodec Token 2023.08
+UniAudio [244] Decoder-only Transformer EnCodec Token 2023.10
+ELLA-V [245] Decoder-only Transformer EnCodec Token 2024.01
+Base TTS [246] Decoder-only Transformer HiFi-GAN+BigVGAN Token 2024.02
+CLaM-TTS [247] Encoder-decoder Transformer BigVGAN Token + MelS 2024.04
+RALL-E [248] Decoder-only Transformer SoundStream Token 2024.05
+ARDiT [249] Decoder-only DiT BigVGAN MelS 2024.06
+VALL-E R [250] Decoder-only Transformer Vocos Token 2024.06
+VALL-E 2 [251] Decoder-only Transformer Vocos Token 2024.06
+Seed-TTS [252] Decoder-only Transformer + DiT Unknown Latent Feature 2024.06
+VoiceCraft [93] Decoder-only Transformer HiFi-GAN Token 2024.06
+XTTS [253] Decoder-only Transformer HiFi-GAN-based Vocoder [253] Token + MelS 2024.06
+CosyVoice [17] Decoder-only Transformer + Flow HiFi-GAN Token 2024.07
+MELLE [254] Decoder-only Transformer HiFi-GAN MelS 2024.07
+VoxInstruct [103] Decoder-only Transformer Vocos Token 2024.08
+Emo-DPO [31] Decoder-only Transformer HiFi-GAN Token + MelS 2024.09
+FireRedTTS [255] Decoder-only Transformer + Flow BigVGAN Token + MelS 2024.09
+CoFi-Speech [256] Decoder-only Transformer BigVGAN Token + MelS 2024.09
+Takin [257] Decoder-only Transformer + Flow HiFi-GAN Token + MelS 2024.09
+HALL-E [258] Decoder-only Transformer EnCodec Token 2024.10
+FishSpeech [259] Decoder-only Transformer Firefly-GAN [259] Token 2024.11
+SLAM-Omni [260] Decoder-only Transformer HiFi-GAN Token + MelS 2024.12
+IST-LM [261] Decoder-only Transformer HiFi-GAN Token + MelS 2024.12
+KALL-E [262] Decoder-only Transformer WaveVAE [262] Latent Feature 2024.12
+IDEA-TTS [263] Transformer Flow-based Vocoder [263] LinS + MelS 2024.12
+FleSpeech [264] Flow-based DiT WaveGAN [141] Latent Feature 2025.01
+Step-Audio [265] Decoder-only Transformer Flow-based Vocoder [265] Token 2025.02
+Vevo [266] Decoder-only Transformer BigVGAN Token + MelS 2025.02
+Spark-TTS [267] Decoder-only Transformer BiCodec [267] Token 2025.03
+Abbreviations: Pit(ch), Ene(rgy), Spe(ed), Pro(sody), Tim(bre), Emo(tion), Env(ironment), Des(cription). MelS: Mel Spectrogram. LinS: Linear Spectrogram.
+IV-B Autoregressive Architectures
+For an autoregressive model in TTS, the probability of the speech frame sequence $\mathbf{y}=(y_{1},y_{2},\dots,y_{T})$ given the input sequence $\mathbf{x}=(x_{1},x_{2},\dots,x_{T})$ can be modeled as:
+$P(\mathbf{y}|\mathbf{x})=\prod_{t=1}^{T}P(y_{t}|y_{<t},\mathbf{x}),$ (2)
+where $y_{t}$ is the predicted output frame at time step $t$, $y_{<t}=(y_{1},\dots,y_{t-1})$ are the previous frames, and $\mathbf{x}$ is the input feature sequence.
+Each frame $y_{t}$ is predicted conditioned on all previous frames and the input sequence $\mathbf{x}$.
+Autoregressive models are powerful in TTS modeling but tend to be slower in generation time compared to non-autoregressive models, making them suitable for applications where quality is prioritized over real-time performance.
+HMM-based Approaches.
+In the realm of controllable TTS, advancements in HMM architectures have significantly enhanced the manipulation of speech elements such as emotion and prosody.
+Yamagishi et al.
+[70] pioneered this field by introducing style-dependent and style-mixed modeling, which allowed precise emulation of human-like emotional nuances and versatile synthesis across various styles by incorporating style as a contextual variable.
+Building on this foundation, Qin et al.
+[268] developed the “average emotion model,” which utilized maximum likelihood linear regression-based adaptation to modulate emotions like happiness and sadness even with limited data, thus advancing the emotional intelligence of synthetic speech systems.
+Furthering expressive variability, Nose et al.
+[117] integrated subjective style intensities and a multiple-regression global variance model into HMMs, addressing over-smoothing and enabling nuanced emotional expressions.
+Lorenzo-Trueba et al.
+[72] expanded on these capabilities with CSMAPLR [72] adaptation, introducing “emotion transplantation” to transfer emotional states between speakers while preserving voice distinctiveness, enhancing personalized human-computer interaction.
+These innovations in HMM architectures have broadened the expressiveness and individuality in synthetic speech, augmenting technological interfaces and paving the way for future developments in adaptive, lifelike speech generation.
 RNN-based Approaches.
-Controllable Text-To-Speech (TTS) technology has seen significant advancements through innovations in neural network architectures, facilitating speech that is both natural-sounding and adaptable in terms of emotion, prosody, and pitch.
-A key breakthrough was the introduction of Tacotron[74], a sequence-to-sequence architecture that effectively integrates prosodic variations, laying the groundwork for precise control over speech attributes.
-Tacotron 2[175] further enhanced this capability by better managing prosodic variability, though it averaged these variations, indicating a need for more sophisticated control methods.
-To address these constraints, Wang et al.[19] introduced Global Style Tokens (GSTs), using an unsupervised approach to encapsulate diverse speech styles into fixed tokens, thus enabling versatile style transfer within the Tacotron framework.
-Skerry-Ryan et al.[201] further advanced this by incorporating prosodic embeddings, providing detailed control over timing and intonation, and significantly improving the replication of emotional expressions in synthetic speech.
+Controllable TTS technology has seen significant advancements through innovations in neural network architectures, particularly RNN-based architectures, enabling natural-sounding speech generation with adjustable emotion, prosody, and pitch.
+Prosody-Tacotron [234] is an extension of the original Tacotron model, designed to improve the prosody (rhythm, intonation, stress patterns) of synthesized speech.
+It builds on the Tacotron framework by introducing additional mechanisms to explicitly control prosodic features, which is a key challenge in TTS systems.
+Wang et al.[19] introduced global style tokens, using an unsupervised approach to encapsulate diverse speech styles into fixed tokens, thus enabling versatile style transfer within the Tacotron framework.
+Similarly, Stanton et al.
+[235] introduce Text-Predicted Global Style Tokens (TP-GST), enhancing Tacotron speech synthesis to predict expressive speaking styles directly from text without auxiliary inputs.
+Experiments show TP-GST generates more prosodic variation and achieves higher listener preference than baselines.
+Skerry-Ryan et al.[234] further advanced this by incorporating prosodic embeddings, providing control over timing and intonation, and significantly improving the replication of emotions in synthetic speech.
 Building on these innovations, emotion-controllable models developed by Li et al.[21] focus on calibrating emotional nuances using emotion embedding networks and style loss alignment, allowing detailed modulation of emotional strength.
@@
 LLM-based Approaches.
-Inspired by the success of large language models (LLMs) in natural language processing (NLP), recent studies have explored leveraging in-context learning for zero-shot TTS generation.
+Inspired by the success of LLMs in natural language processing (NLP), recent studies have explored leveraging in-context learning for zero-shot TTS.
+As shown in Fig.
+3, LLM-based approaches often take the target text or instructions and an optional reference speech clip as input, and utilize autoregressive modeling to generate speech tokens or features, which are then converted into the final waveform by a decoder.
+[…]
 VALL-E [85] is a pioneering work in this area, formulating TTS as a conditional language modeling problem.
@@
 VALL-E R [250] improves phoneme-to-acoustic alignment and adopts codec-merging to boost decoding efficiency and reduce computational overhead.
-VALL-E 2 [212] introduces repetition aware sampling and grouped code modeling for greater stability and faster inference.
-HALL-E [213] adopts a hierarchical post-training framework, effectively managing the trade-off between reducing frame rate and producing high-quality speech.
-Beyond the foundational improvements introduced by VALL-E and its immediate extensions, further advancements have focused on enhancing speech alignment, quality, and robustness.
+VALL-E 2 [251] introduces repetition-aware sampling and grouped code modeling for greater stability and faster inference.
+HALL-E [258] adopts a hierarchical post-training framework, effectively managing the trade-off between reducing the frame rate and producing high-quality speech.
+Beyond the foundational improvements introduced by VALL-E and its extensions, further advancements have focused on enhancing speech alignment, quality, and robustness.
 SpearTTS [237] and Make-a-Voice [239] use semantic tokens to bridge the gap between text and acoustic features.
@@
 CoFi-Speech [256] generates speech in a coarse-to-fine manner via a multi-scale speech coding and generation approach, producing natural and intelligible speech.
-CosyVoice [17] employs supervised semantic tokens to enhance content consistency and speaker similarity in zero-shot voice cloning.
 Similarly, BASE TTS [246] introduces discrete speech representations based on the WavLM [233] self-supervised model, focusing on phonemic and prosodic information.
-SeedTTS [217] also proposes a self-distillation method for speech decomposition and a reinforcement learning approach to enhance the robustness, speaker similarity and controllability of generated speech.
-Based on this framework, Bailing-TTS [218] enhances the alignment of text and speech tokens using a continual semi-supervised learning strategy, enabling high-quality synthesis of Chinese dialect speech.
-Although models using discrete tokens as intermediate representations have achieved notable success in zero-shot TTS, they still face fidelity issues compared to the continuous representation like Mel spectrograms [219, 220].
+SeedTTS [252] also proposes a self-distillation method for speech decomposition and a reinforcement learning approach to enhance the robustness, speaker similarity, and controllability of generated speech.
+Although models using discrete tokens as intermediate representations have achieved notable success in zero-shot TTS, they still face fidelity issues compared to continuous representations like Mel spectrograms [254, 249].
 MELLE [254] optimizes the training objectives and sampling strategy, marking the first exploration of using continuous-valued tokens instead of discrete-valued tokens within the paradigm of autoregressive speech synthesis models.
-Similar to MELLE, ARDiT [220] encodes audio as vector sequence in continuous space and autoregressively generates these sequences by a decoder-only transformer.
-TABLE III: A summary of existing non-autoregressive controllable neural-based methods.
-Method Zero-shot TTS Controlability Model Architectures Acoustic Feature Release Time
-Pit. Ene. Spe. Pro. Tim. Emo. Env. Des. Acoustic Model Vocoder
-FastSpeech[15] Transformer WaveGlow MelS 2019.05
-DWAPI [221] DNN Straight MelS + F0 + Intensity 2020.04
-FastSpeech 2[180] Transformer Parallel WaveGAN MelS 2020.06
-FastPitch[77] Transformer WaveGlow MelS 2020.06
-Parallel Tacotron[222] Transformer + CNN WaveRNN MelS 2020.10
-StyleTagging-TTS [223] Transformer + CNN HiFi-GAN MelS 2021.04
-SC-GlowTTS [224] Transformer + Conv HiFi-GAN MelS 2021.06
-Meta-StyleSpeech [225] Transformer MelGAN MelS 2021.06
-DelightfulTTS[226] Transformer + CNN HiFiNet MelS 2021.11
-YourTTS [82] Transformer HiFi-GAN LinS 2021.12
-DiffGAN-TTS[227] Diffusion + GAN HiFi-GAN MelS 2022.01
-StyleTTS [88] CNN + RNN + GAN HiFi-GAN MelS 2022.05
-GenerSpeech [90] Transformer + Flow-based HiFi-GAN MelS 2022.05
-NaturalSpeech 2 [86] Diffusion Codec Decoder Token 2022.05
-Cauliflow[228] BERT + Flow UP WaveNet MelS 2022.06
-CLONE[181] Transformer + CNN WaveNet MelS + LinS 2022.07
-PromptTTS [101] Transformer HiFi-GAN MelS 2022.11
-Grad-StyleSpeech [182] Score-based Diffusion HiFi-GAN MelS 2022.11
-PromptStyle [229] VITS HiFi-GAN MelS 2023.05
-StyleTTS 2 [89] Diffusion + GAN HifiGAN / iSTFTNet MelS 2023.06
-VoiceBox [189] Flow Matching Diffusion HiFi-GAN MelS 2023.06
-MegaTTS 2 [230] Diffusion + GAN HiFi-GAN MelS 2023.07
-PromptTTS 2 [102] Diffusion Codec Decoder Token 2023.09
-VoiceLDM [231] Diffusion HiFi-GAN MelS 2023.09
-DuIAN-E[232] CNN + RNN HiFi-GAN MelS 2023.09
-PromptTTS++ [233] Transformer + Diffusion BigVGAN MelS 2023.09
-SpeechFlow [190] Flow Matching Diffusion HiFi-GAN MelS 2023.10
-P-Flow [188] Flow Matching HiFi-GAN MelS 2023.10
-E3 TTS [184] Diffusion / Waveform→Unet 2023.11
-HierSpeech++ [191] Hierarchical Conditional VAE BigVGAN MelS 2023.11
-Audiobox [187] Flow Matching EnCodec MelS 2023.12
-FlashSpeech [192] Latent Consistency Model EnCodec Token 2024.04
-NaturalSpeech 3 [87] Diffusion EnCodec Token 2024.04
-InstructTTS [105] Transformer + Diffusion HiFi-GAN Token 2024.05
-ControlSpeech [106] Transformer + Diffusion FACodec Decoder Token 2024.06
-AST-LDM [234] Diffusion HiFi-GAN MelS 2024.06
-SimpleSpeech [186] Transformer Diffusion SQ Decoder Token 2024.06
-DiTTo-TTS [185] DiT BigVGAN Token 2024.06
-E2 TTS [193] Flow Matching Transformer BigVGAN MelS 2024.06
-MobileSpeech [235] ConFormer Decoder Vocos Token 2024.06
-DEX-TTS [183] Diffusion HiFi-GAN MelS 2024.06
-ArtSpeech [236] RNN + CNN HiFI-GAN MelS+Energy+F0+TV 2024.07
-CCSP [237] Diffusion Codec Decoder Token 2024.07
-SimpleSpeech 2 [199] Flow-based Transformer Diffusion SQ Decoder Token 2024.08
-E1 TTS [196] DiT BigVGAN Continuous Token 2024.09
-VoiceGuider [238] Diffusion BigVGAN MelS 2024.09
-StyleTTS-ZS [239] Diffusion + GAN HifiGAN / iSTFTNet Token 2024.09
-NansyTTS [240] Transformer NANSY++ MelS 2024.09
-NanoVoice [241] Diffusion BigVGAN MelS 2024.09
-MS2KU-VTTS [242] Diffusion BigvGAN MelS 2024.10
-MaskGCT [78] Masked Generative Transformers DAC + Vocos Token 2024.10
-Abbreviations: Pit(ch), Ene(rgy), Spe(ed), Pro(sody), Tim(bre), Emo(tion), Env(ironment), Des(cription).
-Timbre involves gender and age.
-MelS and LinS represent Mel Spectrogram and Linear Spectrogram respectively.
-TV represents the vocal tract variables proposed in  [236].
-TABLE IV: A summary of existing autoregressive controllable neural-based methods.
-Method Zero-shot TTS Controlability Model Architectures Acoustic Feature Release Time
-Pit. Ene. Spe. Pro. Tim. Emo. Env. Des. Acoustic Model Vocoder
-Prosody-Tacotron[201] RNN WaveNet MelS 2018.03
-GST-Tacotron[243] CNN + RNN Griffin-Lim LinS 2018.03
-GMVAE-Tacotron[130] CNN + RNN WaveRNN MelS 2018.12
-VAE-Tacotron[18] CNN + RNN WaveNet MelS 2019.02
-DurIAN[244] CNN + RNN MB-WaveRNN MelS 2019.09
-Flowtron[245] CNN + RNN WaveGlow MelS 2020.07
-MsEmoTTS[83] CNN + RNN WaveRNN MelS 2022.01
-VALL-E [85] LLM EnCodec Token 2023.01
-SpearTTS [203] LLM SoundStream Token 2023.02
-VALL-E X [207] LLM EnCodec Token 2023.03
-Make-a-voice [204] LLM BigVGAN Token 2023.05
-TorToise [246] Transformer + DDPM Univnet MelS 2023.05
-MegaTTS [91] LLM + GAN HiFi-GAN MelS 2023.06
-SC VALL-E [205] LLM EnCodec Token 2023.07
-Salle [247] LLM Codec Decoder Token 2023.08
-UniAudio [248] LLM EnCodec Token 2023.10
-ELLA-V [208] LLM EnCodec Token 2024.01
-BaseTTS [216] LLM UnivNet Token 2024.02
-ClaM-TTS [249] LLM BigVGAN MelS+Token 2024.04
-RALL-E [209] LLM SoundStream Token 2024.05
-ARDiT [220] Decoder-only Diffusion Transformer BigVGAN MelS 2024.06
-VALL-E R [211] LLM Vocos Token 2024.06
-VALL-E 2 [212] LLM Vocos Token 2024.06
-Seed-TTS [217] LLM + Diffusion Transformer / Token 2024.06
-VoiceCraft [93] LLM HiFi-GAN Token 2024.06
-XTTS [250] LLM + GAN HiFi-GAN MelS+Token 2024.06
-CosyVoice [17] LLM + Conditional Flow Matching HiFi-GAN Token 2024.07
-MELLE [219] LLM HiFi-GAN MelS 2024.07
-Bailing TTS [218] LLM + Diffusion Transformer / Token 2024.08
-VoxInstruct [103] LLM Vocos Token 2024.08
-Emo-DPO [31] LLM HiFi-GAN Token 2024.09
-FireRedTTS [214] LLM + Conditional Flow Matching BigVGAN-v2 Token 2024.09
-CoFi-Speech [215] LLM BigVGAN Token 2024.09
-Takin [206] LLM HiFi-Codec Token 2024.09
-HALL-E [213] LLM EnCodec Token 2024.10
-Abbreviations: Pit(ch), Ene(rgy), Spe(ed), Pro(sody), Tim(bre), Emo(tion), Env(ironment), Des(cription).
-Timbre involves gender and age.
-MelS and LinS represent Mel Spectrogram and Linear Spectrogram respectively.
-IV-B Control Strategies
-The control strategies in existing controllable TTS can be broadly classified into four categories: style tagging using discrete labels, speech reference prompt for customizing a new speaker’s voice with just a few seconds of voice input, controlling speech style using natural language descriptions, and the instruction-guided mode.
+Similar to MELLE, ARDiT [249] encodes audio as a vector sequence in continuous space and autoregressively generates these sequences by a decoder-only transformer.
+Additionally, some autoregressive methods enable speech editing through natural language instructions.
+VoiceCraft [93] introduces a decoder-only Transformer-based neural codec language model that utilizes causal masking and delayed stacking for token rearrangement.
+This approach allows for bidirectional context-aware speech editing and zero-shot TTS.
+VoiceCraft achieves precise control through text-guided modifications, such as insertion, deletion, and substitution, producing edits that are nearly indistinguishable from unmodified recordings in terms of naturalness.
+InstructSpeech [271] employs a multi-task LLM trained on triplet data (instruction, input, and output speech) with task embeddings and hierarchical adapters.
+It enables fine-grained control over both semantic attributes (content editing) and acoustic properties (emotion, speed) using natural language instructions.
+By leveraging multi-step reasoning, InstructSpeech facilitates free-form editing and efficiently adapts to new tasks.
+IV-C Control Strategies
+The control strategies in existing controllable TTS can be broadly classified into four categories: style tagging using discrete labels or continuous control signals, reference speech prompt for customizing a new speaker’s voice with just a few seconds of voice input, controlling speech style using natural language descriptions, and instruction-guided speech attributes control.
 We illustrate taxonomies of controllable TTS from the perspective of control strategies in Fig.
-3.
+4.
+TABLE V: A summary of control strategies for TTS.
+Control Strategy Control Signal
+Style Tagging Discrete labels, continuous values, latent variables, speech audio
+Speech Reference Prompt Speech audio
+Natural Language Descriptions User input text, speech audio
+Instruction-Guided Control User input text, speech audio
 […]
 Figure 4: A taxonomy of controllable TTS from the perspective of control strategies.
-IV-B1 Style Tagging
-This paradigm typically employs target control attributes, primarily emotion-related controls, as categorical label inputs to enable controllable speech synthesis.
-StyleTagging-TTS [223] utilizes a short phrase or word to represent the style of an utterance and learns the relationship between linguistic embedding and style embedding space by a pre-trained language model.
-However, these methods are limited in expressive diversity, as they can only model a small set of pre-defined styles.
-IV-B2 Reference Speech Prompt
+Style Tagging.
+This paradigm enables controllable speech synthesis by adjusting key attributes such as pitch, energy, speech rate, and emotion.
+These attributes can be controlled using either categorical labels or continuous values.
+In this context, “tagging” refers to the process of assigning a control signal to a specific speech attribute.
+Some approaches utilize discrete labels to control synthesized speech.
+For instance, StyleTagging-TTS[182] employs short phrases or words to represent utterance styles.
+It learns the relationship between linguistic embeddings and style embeddings using a pre-trained language model.
+Emo-DPO[31] enables precise emotion control (e.g., “angry,” “happy”) via Direct Preference Optimization (DPO) [276] with LLM.
+By training on paired emotion-text data, it leverages contrastive learning to distinguish subtle prosodic differences between emotions.
+Users can specify an emotion label to shape the expressive quality of the generated speech.
+Spark-TTS [267] offers both coarse-grained control (e.g., gender, speaking style) and fine-grained adjustments (e.g., precise pitch values, speaking rate).
+Users can modify specially designed tokens through prompts to achieve customized speech synthesis with optional reference speech as input.
+Other methods enable control by adjusting continuous input signals.
+DiffStyleTTS[221] models prosody hierarchically, incorporating coarse-grained implicit style conditions (extracted via GST [19]) and fine-grained explicit features.
+Users can adjust guiding scale factors to control pitch, energy, duration, and prosodic styles.
+DrawSpeech[222] provides an intuitive way to manipulate pitch and energy.
+Users can sketch rough prosody contours (e.g., rising or falling patterns), which the system refines into detailed contours through a sketch-to-contour predictor.
+A diffusion model then generates expressive speech that aligns with the user’s sketched prosody, offering precise control over vocal emphasis and intonation.
+In addition, some approaches regulate speech attributes by modifying latent features rather than direct input signals.
+Cauliflow [186] controls speech rate and pausing by conditioning its flow-based duration model on user-defined parameters ($r_{s}$ for speed, $r_{p}$ for pause frequency) that adjust deviations from dataset averages during synthesis.
+DurIAN-E [195] allows users to manipulate expressive styles.
+It employs variance predictors for prosodic attributes and incorporates Style-Adaptive Instance Normalization (SAIN) layers, which dynamically adjust mel-spectrogram statistics based on predefined style embeddings.
+DiTTo-TTS [205], a diffusion transformer-based TTS model, removes dependencies on domain-specific features such as phonemes and durations while maintaining high performance.
+It controls speech rate by modifying the latent length predicted by a length predictor.
+These methods show great potential in controlling speech attributes by adjusting input signals or latent variables.
+However, these methods are limited in expressive diversity, as they can only model a small set of pre-defined attributes.
+Reference Speech Prompt.
 This paradigm aims to customize a new speaker’s voice with just a few seconds of voice prompt.
@@
 MetaStyleSpeech [184] and StyleTTS [88] use adaptive normalization as a style conditioning method, enabling robust zero-shot performance.
-GenerSpeech [90] introduces a multi-level style adapter to improve zero-shot style transfer for out-of-domain custom voices.
+GenerSpeech [90] introduces a multilevel style adapter to improve zero-shot style transfer for out-of-domain custom voices.
 SC VALL-E [242] facilitates control over synthesized speech’s emotions, speaking styles, and various acoustic features by incorporating style tokens and scale factors.
 ArtSpeech [210] revisits the sound production system by integrating articulatory representations into the TTS framework, improving the physical interpretability of articulation movements.
-To enhance the learning of contextual information and address the challenge of limited voice data from the target speaker, CCSP [237] proposes a contrastive context-speech pretraining framework that learns cross-modal representations combining both contextual text and speech expressions.
+To enhance the learning of contextual information and address the challenge of limited voice data from the target speaker, CCSP [211] proposes a contrastive context-speech pretraining (CCSP) framework that learns cross-modal representations, combining both contextual text and speech expressions.
 DEX-TTS [209] separates styles into time-invariant and time-variant components, enabling the extraction of diverse styles from expressive reference speech.
@@
 ControlSpeech [106] uses bidirectional attention and mask-based parallel decoding to capture codec representations in a discrete decoupling codec space, allowing independent control of timbre, style, and content in a zero-shot manner.
-IV-B3 Natural Language Descriptions
+Natural Language Descriptions.
 Recent studies explore controlling speech style using natural language descriptions that include attributes such as pitch, gender, and emotion, making the process more user-friendly and interpretable.
@@
 NansyTTS [215] combines a TTS trained on the target language with a description control model trained on another language, which shares the same timbre and style representations to enable cross-lingual controllability.
-Considering not all details about voice variability can be described in the text prompt, PromptTTS++ [233] and PromptSpeaker [251] tries to construct text prompts with more details.
+Considering that not all details about voice variability can be described in the text prompt, PromptTTS++ [104] and PromptSpeaker [272] try to construct text prompts with more details.
 PromptTTS 2 [102] designs a variation network to capture voice variability not conveyed by text prompts.
@@
 Audiobox [201] introduces both description-based and example-based prompting, integrating speech and sound generation paradigms to independently control transcript, vocal, and other audio styles during speech generation.
-VoiceLDM [231] and AST-LDM [234] extend AudioLDM [253] to incorporate environmental context in TTS generation by adding a content prompt as a conditional input.
-Building on VoiceLDM, MS2KU-VTTS [242] further expands the dimensions of environmental perception, enhancing the generation of immersive spatial speech.
-IV-B4 Instruction-Guided Control
-The natural language description-based TTS methods discussed above require splitting inputs into content and description prompts, which limits fine-grained control over speech and does not align with other AIGC models.
+VoiceLDM [194] and AST-LDM [203] extend AudioLDM [224] to incorporate environmental context in TTS by adding a content prompt as a conditional input.
+Building on VoiceLDM, MS2KU-VTTS [217] further expands the dimensions of environmental perception, enhancing the generation of immersive speech.
+Instruction-Guided Control.
+The description-based TTS methods discussed above require splitting inputs into content and description prompts, which limits fine-grained control over speech and does not align with other AIGC models.
 VoxInstruct [103] proposes a new paradigm that extends traditional text-to-speech tasks into a general human instruction-to-speech task.
@@
 To enable automatic extraction of the synthesized speech content from raw text instructions, VoxInstruct uses speech semantic tokens as an intermediate representation, bridging the gap in current research by allowing the simultaneous use of both text description prompts and speech prompts for speech generation.
-V Datasets and Evaluation
-TABLE V: A summary of open-source datasets for controllable TTS.
-Dataset Hours #Speakers Labels Lang Release Time
-Pit. Ene. Spe. Age Gen. Emo. Emp. Acc. Top. Des. Env. Dia.
-Taskmaster-1 [254] / / en 2019.09
-Libri-light [255] 60,000 9,722 en 2019.12
-AISHELL-3 [256] 85 218 zh 2020.10
-ESD [257] 29 10 en,zh 2021.05
-GigaSpeech [258] 10,000 / en 2021.06
-WenetSpeech [259] 10,000 / zh 2021.07
-PromptSpeech [101] / / en 2022.11
-DailyTalk [260] 20 2 en 2023.05
-TextrolSpeech [247] 330 1,324 en 2023.08
-VoiceLDM [231] / / en 2023.09
-VccmDataset [106] 330 1,324 en 2024.06
-MSceneSpeech [261] 13 13 zh 2024.07
-SpeechCraft [262] 2,391 3,200 en,zh 2024.08
-Abbreviations: Pit(ch), Ene(rgy)=volume=loudness, Spe(ed)=duration, Gen(der), Emo(tion), Emp(hasis), Acc(ent), Dia(logue), Env(ironment), Des(cription).
-TABLE VI: Common objective and subjective evaluation metrics
+CosyVoice [17] introduces supervised semantic tokens derived from ASR models via vector quantization, enabling precise text-speech alignment.
+It employs an LLM for token generation and flow matching for synthesis.
+Controllability is enhanced through instruction fine-tuning, allowing adjustments in speaker identity, emotion, speaking rate, pitch, and paralinguistic elements (e.g., laughter, breath sounds) via textual instruction.
+Some speech LLMs can also synthesize speech content, but offer less controllability.
+These speech LLMs are agents that take user instructions and call expert TTS models to synthesize speech content.
+AudioGPT [274] is a multimodal LLM designed to process and generate audio data.
+It extends the capabilities of general-purpose LLMs by incorporating audio/speech understanding, synthesis (FastSpeech2 [76]), and style conversion [90] modules.
+SpeechGPT [34] integrates speech and text via discrete representations.
+It employs a three-stage training strategy for cross-modal alignment.
+For speech synthesis control, it uses HiFi-GAN with speaker embeddings and chain-of-modality instructions, enabling instruction-guided speech generation, e.g., one can input “Please read the sentence: Today is a beautiful day.” FunAudioLLM [275] is an LLM enhancing voice interaction between humans and LLMs via SenseVoice [275] (multilingual speech recognition) and CosyVoice [17] (controllable speech synthesis).
+StepAudio [265] introduces a unified 130B speech-text model integrating understanding/generation, with a generative data engine enabling affordable voice cloning and a controllable TTS system (Step-Audio-TTS-3B [265]).
+Its speech synthesis achieves dynamic control through instruction-driven adjustments of dialects, emotions (anger/joy/sadness), singing styles, RAP vocals, and speaking rates.
+V Datasets and Evaluation Metrics
+TABLE VI: A summary of publicly available speech datasets for controllable TTS.
+Dataset Hours (at least) #Speakers (at least) Labels Lang Release Time
+Pit. Ene. Spe. Age Gen. Emo. Emp. Acc. Top. Des. Dia.
+IEMOCAP [277] 12 10 en 2008
+RECOLA [278] 3.8 46 fr 2013
+RAVDESS [279] / 24 en 2018
+CMU-MOSEI [280] 65 1,000 en 2018
+Taskmaster-1 [281] / / en 2019
+AISHELL-3 [282] 85 218 zh 2020
+Common Voice [283] 2,500 50,000 multi 2020
+ESD [284] 29 10 en,zh 2021
+GigaSpeech [285] 10,000 / en 2021
+WenetSpeech [286] 10,000 / zh 2021
+PromptSpeech [101] / / en 2022
+MagicData-RAMC [287] 180 663 zh 2022
+DailyTalk [288] 20 2 en 2023
+TextrolSpeech [243] 330 1,324 en 2023
+CLESC [289] $<$1 / en 2024
+VccmDataset [106] 330 1,324 en 2024
+MSceneSpeech [290] 13 13 zh 2024
+Parler-TTS [273] 50,000 / en 2024
+SpeechCraft [291] 2,391 3,200 en,zh 2024
+Abbreviations: Pit(ch), Ene(rgy)=volume, Spe(ed), Gen(der), Emo(tion), Emp(hasis), Acc(ent), Top(ic), Des(cription), Dia(logue).
+V-A Datasets
+Fully controllable text-to-speech (TTS) systems require large-scale datasets that exhibit extensive diversity and include fine-grained annotations.
+Such datasets provide essential information that enables TTS models to generate highly expressive speech with precise control over various speech attributes.
+In this subsection, we classify speech datasets into three categories based on the type of annotations they provide: (1) tag-based datasets, which include structured attribute labels such as age and gender; (2) description-based datasets, which contain detailed textual descriptions of speech characteristics; and (3) dialogue datasets, which capture natural conversational speech.
+Below, we provide an overview of these datasets and discuss their importance in advancing TTS research.
+Tag-based Datasets.
+Tag-based datasets consist of speech recordings annotated with predefined attribute tags that characterize various aspects of the speech signal [282, 284, 285, 286, 288, 290, 287, 277, 278, 279, 280].
+These attributes include pitch, energy, speed, age, gender, emotion, emphasis, accent, and topic, among others.
+Such datasets serve as a critical resource for developing TTS models with enhanced expressiveness and fine-grained control over speech generation.
+By incorporating attribute labels, models can be trained to adjust specific characteristics dynamically, enabling more personalized and context-aware speech synthesis.
+Description-based Datasets.
+Description-based datasets go beyond structured labels by pairing speech samples with detailed textual descriptions that characterize various speech attributes, such as intonation, prosody, speaking style, and emotional nuances [101, 243, 106, 291, 273].
+Unlike tag-based datasets, which provide predefined categorical labels, description-based datasets enable models to interpret nuanced, free-form textual prompts and generate speech that aligns with complex, context-dependent specifications.
+This category of datasets plays a crucial role in training TTS systems that can respond to natural language descriptions, making them highly suitable for applications requiring fine-grained expressiveness, such as storytelling, audiobooks, and personalized voice assistants.
+Dialogue Datasets.
+Dialogue datasets [281, 288, 287] capture multi-turn conversational speech between two or more speakers, focusing on the natural flow of human interaction, including turn-taking, contextual dependencies, and prosodic variations.
+These datasets are particularly valuable for training conversational TTS models that can generate dynamic and contextually appropriate speech for chatbots and interactive applications.
+By incorporating dialogue-specific characteristics such as speaker intent, pauses, and conversational nuances, models trained on these datasets can produce more natural and engaging interactions for real-world scenarios.
+By leveraging these different categories of speech datasets, researchers can develop more advanced and flexible TTS models capable of generating speech that is not only intelligible but also expressive, context-sensitive, and highly controllable.
+We summarize publicly available datasets in Table VI.
+V-B Evaluation
+TABLE VII: Common objective and subjective evaluation metrics.
 Metric Type Eval Target GT Required
-MCD [263] Objective Acoustic similarity
-PESQ [264] Objective Perceptual quality
-WER [265] Objective Intelligibility
-MOS [266] Subjective Preference
-CMOS [267] Subjective Preference
-GT: Ground truth.
-V-A Datasets
-Achieving fully controllable TTS requires large-scale datasets with rich diversity and fine-grained annotations.
-In this subsection, we categorize speech datasets into four types according to the labels they provide, i.e., attribute tags such as age and gender, description, environment, and dialogues.
-Tag-based speech datasets.
-Tag-based datasets [255, 256, 257, 258, 259, 260, 261] are specialized collections of speech data that include metadata tags, such as pitch, emotion, and energy, age, gender, and so on, to enhance the expressiveness and control of TTS systems.
-These datasets not only provide audio-text pairs but also offer additional labels that guide the model in generating diverse and context-aware speech outputs.
-Description-based speech datasets.
-Description-based TTS datasets [101, 247, 231, 106, 262] contain speech data paired with detailed textual descriptions of the speech attributes or characteristics, such as emotion, speed, intonation, and style.
-These datasets enable training TTS models that interpret descriptive prompts to generate expressive and context-aware speech.
-Speech environment datasets.
-Speech environment datasets [231, 268] are collections of speech data annotated with environmental labels, such as park, stadium, office, or street, capturing the acoustic characteristics of various real-world settings.
-These datasets are crucial for training models to handle diverse acoustic scenarios, improving their robustness and adaptability.
-Dialogue speech datasets.
-Dialogue speech datasets [260, 254] capture conversational speech between two or more people, focusing on the natural flow, turn-taking, and context of human dialogue.
-These datasets are crucial for training systems in applications like chatbots, virtual assistants, and interactive storytelling.
-We summarize open-source speech datasets for controllable TTS in Table V.
-V-B Evaluation
+MCD [292$\downarrow$ Objective Acoustic similarity
+FDSD [293$\downarrow$ Objective Acoustic similarity
+WER [294$\downarrow$ Objective Intelligibility
+Cosine [295, 296$\downarrow$ Objective Speaker similarity
+PESQ [297$\uparrow$ Objective Perceptual quality
+SNR [298$\uparrow$ Objective Perceptual quality
+MOS [299$\uparrow$ Subjective Preference
+CMOS [300$\uparrow$ Subjective Preference
+AB Test Subjective Preference
+ABX Test Subjective Perceptual similarity
+GT: Ground truth, $\downarrow$: Lower is better, $\uparrow$: Higher is better.
 The performance of controllable TTS often requires objective and subjective evaluation. We introduce common evaluation metrics in this subsection.
@@
 Mel Cepstral Distortion (MCD) [292] measures the spectral distance between synthesized and reference speech, reflecting how closely the generated audio matches the target in terms of acoustic features.
-For intelligibility, the Word Error Rate (WER) [265] is used, comparing transcriptions of synthesized speech to the input text via automated speech recognition.
+A lower MCD value indicates a higher similarity between synthesized and reference speech, meaning better speech synthesis quality.
+Typically, an MCD value below 4 suggests good quality, while values above 6 may indicate significant distortion.
+The MCD is computed as follows:
+$MCD=\frac{10}{\ln 10}\cdot\sqrt{2\sum_{d=1}^{D}(c_{d}^{(syn)}-c_{d}^{(ref)})^{2}},$ (3)
+where $c_{d}^{(syn)}$ represents the d-th Mel Cepstral Coefficient (MCC) of the synthesized speech, $c_{d}^{(ref)}$ represents the d-th MCC of the reference speech, $D$ is the number of MCC, and $\frac{10}{\ln 10}\approx 4.342$ is a constant factor that converts the logarithm to a decibel scale.
+Fréchet DeepSpeech Distance (FDSD) [293] is another metric designed to evaluate the quality and naturalness of synthesized speech.
+It is inspired by the Fréchet Inception Distance (FID) [301] used in image generation but adapted to speech by leveraging a deep speech recognition model.
+FDSD measures the statistical distance between the distributions of real (reference) and synthesized speech in the feature space of a pretrained speech recognition model, such as Deep Speech [302].
+By comparing the mean and covariance of the extracted feature representations, FDSD provides a perceptually relevant assessment of speech synthesis quality.
+A lower FDSD means the synthesized speech is more similar to real speech.
+FDSD can be computed as:
+$FDSD=||\mu_{s}-\mu_{r}||^{2}+\text{Tr}(\Sigma_{s}+\Sigma_{r}-2(\Sigma_{s}\Sigma_{r})^{1/2}),$ (4)
+where $\mu_{s}$ and $\Sigma_{s}$ are the mean and covariance of the embeddings from the synthesized speech, $\mu_{r}$ and $\Sigma_{r}$ are the mean and covariance of the embeddings from the real (reference) speech, $||\mu_{s}-\mu_{r}||^{2}$ represents the squared Euclidean distance between the means, $\text{Tr}(\cdot)$ denotes the trace of a matrix, and $(\Sigma_{s}\Sigma_{r})^{1/2}$ is the geometric mean of the covariance matrices.
+For intelligibility, the Word Error Rate (WER) [294] is used.
+It measures the difference between the recognized transcript and the reference transcript by computing the number of errors made in the transcription process.
+WER is computed as:
+$WER=\frac{S+D+I}{N},$ (5)
+where $S$ is the number of substitutions (wrong word in place of the correct word), $D$ is the number of deletions (missed words), $I$ is the number of insertions (extra words added), and $N$ is the total number of words in the reference transcript.
+Cosine similarity (on speaker embeddings) measures similarity between speaker embeddings of synthesized and reference speech.
+It can be used to evaluate zero-shot TTS (voice cloning) methods, where higher values indicate better speaker similarity.
+Given two speaker embeddings, $\mathbf{e_{1}}$ and $\mathbf{e_{2}}$, their cosine similarity is defined as:
+$CosSim(\mathbf{e_{1}},\mathbf{e_{2}})=\frac{\mathbf{e_{1}}\cdot\mathbf{e_{2}}}{\|\mathbf{e_{1}}\|\|\mathbf{e_{2}}\|},$ (6)
+where speaker embeddings can be extracted from a pre-trained speaker embedding model (e.g., ECAPA-TDNN [295] and x-vectors [296]).
 Perceptual Evaluation of Speech Quality (PESQ) [297] is another objective metric designed to evaluate speech quality by comparing degraded audio with a clean reference.
-It is widely used in telecommunications and speech synthesis, PESQ models human auditory perception, producing a score (typically 1–4.5) that reflects intelligibility and distortion under various conditions, including noise or compression.
+It is widely used in telecommunications and speech synthesis.
+PESQ models human auditory perception, producing a score in the range $[-0.5,-4.5]$ that reflects intelligibility and distortion under various conditions, including noise or compression.
+PESQ involves complex perceptual modeling, its core components can be summarized as:
+$PESQ=a_{0}+a_{1}\cdot D_{frame}+a_{2}\cdot D_{time},$ (7)
+where $D_{frame}$ is the frame-by-frame perceptual distortion, $D_{time}$ is the time-domain distortion, and $a_{0},a_{1},a_{2}$ are regression coefficients. One can refer to [297] for details.
+Signal-to-Noise Ratio (SNR) measures the ratio of signal power to noise power.
+A higher SNR indicates a cleaner signal with less noise, while a lower SNR suggests that noise is dominating the signal.
+However, in TTS, noise can come from different sources, such as artifacts from vocoders, neural network distortions, or background noise in dataset recordings.
+A direct computation of SNR in TTS requires a reference clean speech signal ($x[n]$), a synthesized (or noisy) speech signal ($y[n]$), and extracting the noise component ($e[n]=y[n]-x[n]$) from the synthesized signal.
+The SNR for TTS systems can be computed as:
+$SNR=10\log_{10}\left(\frac{P_{\text{signal}}}{P_{\text{noise}}}\right),$ (8)
+where $P_{\text{signal}}=\frac{1}{N}\sum_{n=1}^{N}x[n]^{2}$ and $P_{\text{noise}}=\frac{1}{N}\sum_{n=1}^{N}e[n]^{2}$.
 Subjective Evaluation Metrics.
 The Mean Opinion Score (MOS) [299] is the most commonly used subjective metric.
-In MOS evaluations, listeners rate the naturalness of synthesized speech on a scale (e.g., 1 to 5), where higher scores indicate better quality.
-MOS captures human perception effectively but requires substantial resources for large-scale evaluations.
+In MOS evaluations, listeners rate various aspects, such as naturalness, expressiveness, quality, intelligibility, et al., of synthesized speech on a scale from 1 to 5, where higher scores indicate better quality.
+MOS captures human perception effectively but is expensive for large-scale evaluations.
 Comparison Mean Opinion Score (CMOS) [300] further evaluates relative quality differences between two TTS audio samples.
@@
 CMOS is used to measure subtle improvements in TTS systems, complementing absolute MOS ratings.
-Table VI summarizes common evaluation metrics for TTS.
+MOS and CMOS scores are computed as the average scores across all listeners:
+$MOS/CMOS=\frac{1}{N}\sum_{i=1}^{N}s_{i},$ (9)
+where $s_{i}$ is the score given by the $i$-th listener, and $N$ is the number of listeners.
+AB and ABX tests are also popular in evaluating TTS methods.
+An AB test involves presenting two versions of a synthesized speech (from different TTS models) to human listeners and asking them to choose which they prefer.
+The goal is to assess which model produces better-sounding speech based on certain criteria, such as naturalness, intelligibility, or clarity.
+In an ABX test, listeners compare two synthesized speech samples to a reference speech sample and determine which one is closer in terms of timbre, prosody, emotion, and other relevant features.
+ABX tests are widely used in evaluating zero-shot TTS methods.
+The AB/ABX test score for a model $m$ is:
+$Score_{AB}/Score_{ABX}=\frac{N_{m}}{N},$ (10)
+where $N_{m}$ represents the number of listeners who prefer the speech synthesized by model $m$, and $N$ denotes the total number of listeners.
+Table VII summarizes widely used metrics for TTS.
 VI Challenges and Future Directions
@@
 In this section, we delve into the primary challenges and analyze their underlying reasons.
-Controllable granularity.
+Controllability.
 A critical challenge in controllable TTS is determining what aspects of speech should be controlled and how to control speech characteristics at a specific granularity.
@@
 Although some LLM-based TTS methods such as VoxInstruct [103] can control various aspects of speech through attribute descriptions, determining the appropriate level of granularity for control and devising methods to achieve precise control at a specific granularity or to enable multiscale and fine-grained control remains a significant challenge.
-Feature extraction and representation.
+Feature Disentanglement and Representation.
 Achieving fully controllable TTS needs good feature disentanglement.
@@
 High-quality, diverse, and appropriately annotated datasets are essential for training controllable TTS systems.
-However, such datasets are scarce and difficult to construct.
-To achieve controllable TTS, training data must encompass a wide range of styles, emotions, accents, and prosodic variations to enable versatile control because limited diversity in datasets can restrict the model’s ability to generalize across unseen styles or emotions.
-Although there are some TTS datasets, such as LibriTTS [272], Gigaspeech [258], and TextrolSpeech [247], their diversity is still not enough for fully controllable TTS due to the lack of corpus of diverse content such as comedies, thrillers, cartoons, etc.
-Constructing large-scale datasets with rich diversity is also expensive and time-consuming.
+However, such datasets are scarce and costly to construct.
+In addition, training data must encompass a wide range of styles, emotions, accents, and prosodic variations to enable versatile control because limited diversity in datasets can restrict the model’s ability to generalize across unseen styles or emotions.
+Although there are some large-scale datasets, such as LibriTTS [306], Gigaspeech [285], and TextrolSpeech [243], their diversity is still not enough for fully controllable TTS due to the lack of corpora for diverse content and scenarios such as comedies, thrillers, and cartoons.
+Constructing these large-scale datasets with rich diversity is expensive and time-consuming.
 Another obstacle is that creating datasets with fine-grained, attribute-specific annotations is labor-intensive and costly.
-Besides, manual annotation of speech attributes requires expert knowledge and is prone to inconsistencies and errors, particularly for subjective qualities like emotion.
+In addition, manual annotation of speech attributes requires expert knowledge and is prone to inconsistencies and errors, particularly for subjective attributes like emotion.
 Currently, most datasets provide only coarse labels, such as gender, age, or a limited range of emotions.
-While some datasets, such as SpeechCraft [262], include natural language descriptions of speech attributes, no existing dataset offers fine-grained variations and annotations within the speech of the same speakers.
-Available datasets for controllable TTS are summarized in Table V.
+While some datasets, such as SpeechCraft [291] and Parler-TTS [273], include natural language descriptions of speech attributes, no existing dataset offers fine-grained variations and annotations within the speech of the same speakers.
+Publicly available datasets for controllable TTS are summarized in Table VI.
 Generalization Ability.
-The ability of a TTS system to generalize effectively is crucial for producing natural, high-quality speech across a wide range of conditions, such as unseen speakers, languages, or topics.
+The ability of a TTS system to generalize effectively is crucial for producing natural, high-quality speech across a wide range of conditions, such as unseen speakers, languages, and topics.
 However, achieving robust generalization remains a significant challenge for modern TTS methods due to various factors.
-Zero-shot controllable TTS [78, 92] aims to synthesize speech for unseen speakers with various speech customization such as emotion using minimal reference audio, which can offer flexibility for personalized voice generation.
-However, it faces significant challenges, including capturing unique speaker characteristics from limited data, accurately reproducing prosody and style, and disentangling speaker identity from other audio attributes like emotion or noise.
-Multilingual generalization [250, 273] in TTS refers to the ability to synthesize natural and intelligible speech across multiple languages, including those not seen during training.
+Zero-shot controllable TTS [78, 92] aims to synthesize speech for unseen speakers with various speech customizations, such as emotion, using minimal reference audio, which can offer flexibility for personalized voice generation.
+However, it faces significant challenges, including capturing unique speaker characteristics from limited data, accurately reproducing prosody and style, and disentangling speaker identity from other attributes such as emotion and background noise.
+Multilingual & low-resource generalization [253, 307] in TTS refers to the ability to synthesize natural and intelligible speech across multiple languages, including those not seen during training.
 This capability is essential for applications like cross-lingual communication, multilingual virtual assistants, and speech synthesis for low-resource languages [308].
-Multilingual generalization still faces many challenges such as linguistic diversity and mismatch and the scarcity of data.
+Multilingual generalization still faces many challenges, such as linguistic diversity, mismatch, and the scarcity of data.
 Cross-lingual speaker generalization is another hurdle, as preserving speaker identity across languages can lead to artifacts.
-Domain adaptation [275] in TTS refers to tailoring a pre-trained TTS model to generate speech for a specific domain or context, such as medical terminology and conversational speech.
+Domain adaptation [309] in TTS refers to tailoring a pre-trained TTS model to generate speech for a specific domain or context, such as medical terminology and debate.
 One challenge is that many specialized domains lack sufficient high-quality annotated data for fine-tuning.
-Besides, adapting prosody, intonation, and speaking style to match domain-specific requirements such as comic dialogue is complex.
+In addition, adapting prosody, intonation, and speaking style to match domain-specific requirements such as comic dialogue is complex.
 Failing to capture domain-specific nuances can make speech sound unnatural or inconsistent with the target context.
@@
 High latency is a major issue, as existing controllable TTS models [78, 101, 102, 103] often necessitate autoregressive processes to synthesize speech.
-The inference time of these models ranges from several to tens of seconds.
-This can be especially problematic for real-time applications like live broadcasting or interactive systems.
-Additionally, the challenge of balancing granularity and efficiency arises, as finer controls demand higher-resolution data and more precise models, leading to increased resource requirements and inefficient training and inference.
+The inference time for these models can range from several seconds to tens of seconds for a short audio clip (e.g., 5 to 10 seconds), and they do not support streaming synthesis.
+This can be particularly problematic for real-time applications like live broadcasting or interactive systems.
+Additionally, balancing granularity and efficiency presents a challenge.
+Finer controls require higher-resolution data and more precise modeling, which increases resource demands and results in inefficient training and inference.
 Another major obstacle lies in the trade-off between model complexity and performance.
-State-of-the-art controllable TTS systems often rely on large neural networks such as LLMs with billions of parameters, which provide superior naturalness and expressiveness but demand significant computational resources.
+SOTA controllable TTS systems often rely on LLMs with billions of parameters, which provide superior naturalness and expressiveness but demand significant computational resources.
 Simplifying these architectures can lead to quality degradation, including artifacts, unnatural prosody, or limited expressiveness.
@@
 […]
-Figure 4: A summary of current major challenges for fully controllable TTS.
+§$\S$§VI-A
+Figure 5: A summary of current challenges for fully controllable TTS.
 VI-B Future Directions
+Figure 6: Future directions and their top-3 challenges.
 In this survey, we conduct a comprehensive investigation and analysis of existing TTS methods, particularly on controllable TTS technologies.
-While these methods show great potential in real-world applications, there are still some limitations that need to be addressed.
+While these methods show great potential in applications, numerous limitations and challenges still need to be addressed.
+Hence, we are still in the early stage of controllable speech synthesis.
 Based on our observations, we outline several promising future directions as follows:
-Fine-grained speech synthesis by natural language description.
-Using natural language description to synthesize human speech with fine-grained control over various audio attributes is currently underexplored.
+Instruction-Guided Fine-grained Speech Synthesis.
+Using natural language description to synthesize human speech with fine-grained control over various speech attributes is currently underexplored.
 Most of the existing works can only control a fixed number of attributes of the synthesized speech.
-Although a few works show great control of emotion, timbres, pitch, gender, and styles, e.g., VoxInstruct [103] and CosyVoice [17], they can frequently synthesize unwanted speech clips.
-Users need to synthesize multiple times to get satisfactory speech.
-Fine-grained speech editing by natural language description.
+Although a few works show great control of emotion, timbres, pitch, gender, and styles, e.g., VoxInstruct [103] and CosyVoice [17], they can frequently synthesize unwanted speech clips inconsistent with user instructions.
+Users often need to synthesize multiple times to get satisfactory speech.
+Instruction-Guided Fine-grained Speech Editing.
 Speech or audio editing has been studied for a long time.
 However, existing methods usually train conditional models and adjust a fixed number of conditional inputs to modify the attributes of synthesized speech, thus lacking fine-grained manipulations [95, 94].
-Therefore, how to learn disentangled speech representations for speech attributes while supporting editing by using natural language description is worthy of investigation.
+Therefore, how to learn disentangled speech representations for speech attributes while supporting instruction-guided editing is worthy of investigation.
 Expressive Multi-modal Speech Synthesis.
-Synthesizing speech from multi-modal data such as texts, images, and videos is an appealing research topic due to its various applications in the industry such as storytelling, filming, and gaming.
-Although there are several related works on this task [6, 24, 276, 277], few of them can fully extract useful information from multi-modal data.
+Synthesizing speech from multi-modal data such as texts, images, and videos is an appealing research topic due to its various applications in the industry, such as storytelling, filming, and gaming.
+Although there are several related works on this task [6, 24, 310, 311], they are still limited in extracting the required information from multimodal data.
 Particularly, synthesizing engaging speech and expressive voiceover for complex visual content sees great opportunities in the future.
 Natural and Emotional Conversational TTS.
-Speech conversational TTS have come out for several decades but remained as cascaded systems for a long time and cannot generate natural and emotional speech.
+Conversational TTS has been studied for several decades, but remained as cascaded systems for a long time, limiting its ability to generate natural and expressive speech.
 These systems are not context-aware, making the synthesized speech sound robotic.
-With the advent of LLMs, existing TTS technologies were directly introduced by simply synthesizing speech from the text generated by LLMs [33].
-However, context-aware conversational TTS with rich emotion and good naturalness has not been well studied.
+With the advent of LLMs, existing TTS technologies are directly introduced to synthesize speech using discrete speech tokens [33, 34].
+However, context-aware conversational TTS with rich emotion and naturalness has not been well studied.
 Zero-shot Long Speech Synthesis with Emotion Consistency.
-Zero-shot TTS emerged in recent years to achieve voice cloning and speech style imitation without fine-tuning, making them more practical in real scenarios [78, 194, 17].
-However, synthesizing long speech with rich emotion and style variation in a zero-shot setting remains challenging due to the lack of rich speech information in short reference audio clips.
-Addressing this issue will make a big step towards fully controllable zero-shot TTS.
-Efficient TTS by natural language description.
-Synthesizing speech with natural language description usually involves training large language encoders and bridge nets between the two modalities which can bring about much more computation overhead compared to previous TTS methods.
-The inference time is also relatively slow, e.g., existing methods usually take tens of seconds to synthesize a short speech audio clip of less than 10 seconds [17, 104].
-Therefore, efficient text and speech modeling and interaction is critical for natural language description-based TTS systems.
-VII Discussion
-VII-A Impacts of Controllable TTS
-Applications.
-Controllable TTS systems allow fine-grained manipulation of speech attributes such as pitch, emotion, speaking rate, and style, enabling a wide range of applications across industries.
-One major application is in virtual assistants and customer support systems, where controllable TTS ensures tailored and context-aware responses.
-For instance, a virtual assistant can speak in a calm tone for technical support but switch to an enthusiastic tone when presenting promotional offers, enhancing user experience.
-In the entertainment industry, controllable TTS is invaluable for creating dynamic voiceovers, audiobooks, and gaming characters.
-It enables precise adjustments in tone and delivery, allowing audiobooks to reflect character emotions and gaming characters to exhibit personality traits that align with their roles.
-Similarly, in education, TTS systems can adapt speaking styles to suit different learners, such as adopting a slow, clear tone for language learning or an engaging, storytelling style for children’s content.
-Controllable TTS is also transformative in assistive technologies, where it can generate speech that reflects the user’s intended emotion or personality.
-This is particularly impactful for individuals with speech impairments, enabling more expressive and natural communication.
-In content localization, controllable TTS systems adjust speech characteristics to suit regional and cultural preferences, ensuring a natural fit for global audiences.
-Additionally, it finds applications in human-computer interaction research, enabling adaptive dialogue systems that modify speech dynamically based on user mood or environment.
-By offering this flexibility, controllable TTS systems enhance accessibility, personalization, and engagement across various domains.
-Deepfakes.
+Zero-shot TTS is capable of voice cloning and speech style imitation without fine-tuning, making it practical in real scenarios [78, 229, 17].
+However, synthesizing long speech with rich emotion and style variation in a zero-shot setting remains challenging due to the limited information in short reference audio clips.
+Addressing this issue will make a big step towards personalized zero-shot TTS.
+Instruction-Guided Efficient TTS.
+Synthesizing speech with user instructions usually involves training language model-based codecs and bridge nets between different modalities, leading to much more computation overhead than previous TTS methods.
+The inference time is also relatively long, e.g., existing instruction-guided methods usually take tens of seconds to synthesize a short speech of less than 10 seconds [17, 104].
+Therefore, efficient text and speech modeling is critical for instruction-guided TTS systems.
+VII Impacts of Controllable TTS
+VII-A Applications
+Controllable TTS enables fine-grained manipulation of speech attributes like pitch, emotion, and style, making it valuable across industries.
+For example, in virtual assistants and customer support, controllable TTS ensures context-aware responses, such as a calm tone for technical help or an enthusiastic pitch for promotions.
+In entertainment, it enhances voiceovers, audiobooks, and gaming characters by adjusting tone and delivery to match emotions and personalities.
+Education benefits from adaptive TTS, like slow, clear articulation for language learning or engaging narration for children.
+For assistive technologies, controllable TTS empowers individuals with speech impairments to express emotions naturally.
+In content localization, it adapts speech to cultural preferences, ensuring a seamless experience for global audiences.
+Additionally, in human-computer interaction, it enables adaptive dialogue systems that adjust speech based on user mood or environment.
+By offering flexibility and expressiveness, controllable TTS enhances accessibility, personalization, and engagement across diverse applications.
+VII-B Security Issues
+One major security issue brought by controllable TTS is deepfakes.
 A deepfake is a type of synthetic media in which a person in an existing image, video, or audio recording is replaced with someone else’s likeness or voice.
-This technology uses deep learning, particularly GANs [278], to create highly realistic, but fabricated, content.
-While deepfakes are most commonly associated with video manipulation, such as face swapping [279], they can also apply to audio, enabling the creation of synthetic speech that mimics a specific person’s voice, which is well known as voice cloning.
-Voice cloning, especially few-shot [280] and zero-shot TTS [85, 78], poses a significant threat to systems that rely on voice authentication, such as banking, customer service, and other identity verification processes.
-With a convincing clone of someone’s voice, attackers could potentially impersonate individuals to gain unauthorized access to sensitive information or accounts.
+This technology uses deep learning, particularly GANs [312], to create highly realistic but fabricated content.
+While deepfakes are most commonly associated with video manipulation, such as face swapping [313], they can also be applied to audio, enabling the creation of synthetic speech that mimics a specific person’s voice, which is well known as voice cloning.
+Voice cloning, especially few-shot [314] and zero-shot TTS [85, 78], poses a significant threat to systems that rely on voice authentication, such as banking, customer service, and other identity verification processes, allowing attackers to impersonate individuals to gain unauthorized access to sensitive information or accounts.
+Another issue is adversarial attacks [315].
+Attackers can manipulate pitch, prosody, and phoneme duration to fool automatic speech recognition and speaker verification.
+It also poses risks in data poisoning [316], where adversarially modified samples degrade model performance.
+On the other hand, controllable TTS can strengthen adversarial defenses by generating diverse speech variations for adversarial training and counter-adversarial augmentation [317].
 To address these concerns, it’s essential to establish robust security protocols, consent-based regulations, and public awareness around voice cloning.
 Furthermore, advancements in detecting voice clones are equally important to help distinguish genuine voices from synthesized ones, protecting both individuals and organizations from potential misuse.
-VII-B Limitation of this Survey
-Although we conduct a comprehensive survey of controllable TTS in this paper, there are some limitations we want to address in the future: 1) A unified benchmark method will be provided to evaluate controllable TTS methods.
-2) Detailed analysis and control strategies of each specific speech attribute will be provided in an updated version of this paper.
-3) The methodologies for feature disentanglement are crucial for controllable TTS but are not adequately discussed.
 VIII Conclusion
-In this survey paper, we first elaborate on the general pipeline for controllable TTS, followed by a glimpse of uncontrollable TTS methods.
-Then we comprehensively review existing controllable methods from the perspectives of model architectures and control strategies.
-Popular datasets and commonly used evaluation metrics for controllable TTS are also summarized in this paper.
-Besides, the current challenges are deeply analyzed and the promising future directions are also pointed out.
+In this survey, we have first elaborated on the general pipeline for controllable TTS, followed by a glimpse of uncontrollable TTS methods from the perspective of local and global speech modeling.
+Then, we have comprehensively reviewed existing controllable methods from the perspectives of model architectures and control strategies.
+Popular datasets and commonly used evaluation metrics for controllable TTS were also summarized in this paper.
+Besides, the current challenges were deeply analyzed, and the promising future directions were also pointed out.
 To the best of our knowledge, this is the first comprehensive survey for controllable TTS.
-Writing a comprehensive survey is a challenging task and an iterative process.
-Hence, we will regularly update this survey to offer researchers and practitioners a continuously evolving resource for understanding controllable speech synthesis.